### v1

# Distinction of the Steinberg representation with respect to split symmetric subgroups

Abstract
We study the distinction of the Steinberg representation of a split reductive group $G$ with respect to a split symmetric subgroup $H\subset G$. We do that both over a $p$ adic field and over a finite field. We relate these distinction problems to problems about determining the existence of a non zero harmonic function on certain hyper graphs related to $X=G/H$. Under these assumptions we verify the relative local Langlands conjecture for the Steinberg representation by showing that over a $p$ adic field the Steinberg representation is $H$-distinguished if and only if its Langalnds parameter factors through the dual group of $X$.
1 Introduction
Let $F$ be a field which is either a non archimedean local field of residual characteristic not 2, or a finite field of characteristic not 2.
Let $\mathbf{G}$ be a connected reductive group split over $F$ and let $G=\mathbf{G}(F)$. Let $\sigma:\mathbf{G}\rightarrow\mathbf{G}$ be an algebraic involution defined over $F$. We denote by $\mathbf{H}$ the fixed points of $\sigma$, and by $H$ the $F$-points of $\mathbf{H}$. We call $X=G/H$ a symmetric space.
A fundamental problem in relative representation theory is the study of $H$-distinguished representations of $G$, that is, smooth representations $\pi$ of $G$ such that $Hom_{H}(\pi,\mathbb{C})\neq 0$. A special representation of $G$ is the Steinberg representation [Sha70]. In this paper, we study the distinction of the Steinberg representation with respect to the symmetric subgroup $H$.
The symmetric space $X$ is called quasi-split if there is a Borel subgroup $B\subset G$ such that $B\cap\sigma(B)$ is a torus. Our first result is that $X$ being quasi split is a necessary condition for $St$ being $H$-distinguished.
In the case of $F$ being a finite field we prove that under the assumption of $H$ being split over $F$, this is a also a sufficient condition for the distinction of the Steinberg representation. In the case of $F$ being a non archimedean local field it is not a sufficient condition. In the example given by $G=GL_{2n+1}(F)$ and $H=GL_{n}(F)\times GL_{n+1}(F)$, $X=G/H$ is quasi split but in this case $St$ isn’t $H$ distinguished. We prove that this example is essentially the unique example for which this happens, see corollary 7.9.
1.1 Main results
We discuss both the case of the Steinberg rpersentation over a finite filed and over a non archimedean local field.
In the case where $F$ is a finite field, there is a unique irreducible representation of $G$ called the Steinberg representation [Ste51] and denoted by $St$. In the case where $F$ is a local field, for any unramified character $\chi:G\rightarrow\mathbb{C}^{\times}$ there is a corresponding Steinberg representation $St_{\chi}$ [Sha70].
An important notion which will appear repeatedly in this work is the notion of $X$ being quasi-split.
Definition 1.1.
We say that $X$ is quasi split if $G$ has a Borel subgroup $B$ such that $\sigma(B)\cap B$ is a maximal torus of $B$.
In the case of a finite field we prove the following:
Theorem 1.2.
Let $F$ be a finite field. Under the assumption that both $G$ and $H$ are split over $F$, $St$ is $H$ distinguished if and only if $X$ is quasi split.
Now assume that $F$ is a non archimedean local field. Let $\mathcal{O}$ be integer ring of $F$. We will assume that both $G$ and $\sigma$ aredefined over $\mathcal{O}$.
In this case, $X$ being quasi split isn’t a sufficient condition, as can be seen from the following proposition.
Proposition 1.3.
Let $F$ be a non archimedean local field. Let $G=GL_{2n+1}(F)$, $H=GL_{n}(F)\times GL_{n+1}(F)$ and $X=G/H$, Then:
  1. 1.
    $X$ is quasi split.
  2. 2.
    For any unramified character $\chi:G\rightarrow\mathbb{C}^{\times}$, the $G$ representation $St_{\chi}$ isn’t $H$-distinguished.
Let $\mathbf{Z}=Z(\mathbf{G})$ be the center of $\mathbf{G}$. We write $\mathbf{G}/\mathbf{Z}$ as a product of adjoint simple groups $G/Z=G_{1}\times...\times G_{m}$. The involution $\sigma$ acts on $\mathbf{G}/\mathbf{Z}$.
Let $T$ be a split torus of $G$, let $T^{0}$ be the maximal compact subgroup of $T$ and let $W_{aff}=N_{G}(T)/T^{0}$ be the affine Weyl group. Let $l$ be the natural length function on $W_{aff}$, denote by $\Omega=l^{-1}(0)$. Let $\omega:G\rightarrow\Omega$ be the natural projection, given by action on the $\Omega$ colors of chambers in the extended Bruhat Tits building of $G$.
Let $\mathcal{C}$ be a chamber in the extended Bruhat Tits buidling of $G$ that is $\sigma$ stable. $\Omega$ acts on $\mathcal{C}$ and $\sigma$ also acts on $\mathcal{C}$. Let $\Omega_{H,\mathcal{C}}=\{\omega(h)|h\in H,h\mathcal{C}=\mathcal{C}\}$ be a subgroup of $\Omega$, we will prove that $\Omega_{H,\mathcal{C}}\subset\Omega$ doesn’t depend on $\mathcal{C}$ (see proposition 8.3). We denote $\Omega_{H}=\Omega_{H,\mathcal{C}}$ for some choice of $\mathcal{C}$.
The following character was introduced in the context of the distinction of the Steinberg representation by Broussous in [Bro14].
Definition 1.4.
Let $\chi_{0}:G\rightarrow\mathbb{C}^{\times}$ be the quadratic unramified character defined by $\chi_{0}(g)=sgn(\omega(g))$, where $sgn$ is the sign of the permutation of $\omega(g)$ on the vertices of the fundamental alcove.
Theorem 1.5.
Let $F$ be a non archimedean local field, assume that $G$ and $H$ are split over $F$. Let $\chi:G\rightarrow\mathbb{C}^{\times}$ be an unramified character, and let $St_{\chi}$ be the corresponding Steinberg representation. Then, $St_{\chi}$ is $H$-distinguished if all the following conditions hold:
  1. 1.
    $X$ is quasi-split.
  2. 2.
    Write $\mathbf{G}/\mathbf{Z}=G_{1}\times...\times G_{m}$. There is no $\sigma$ invariant $G_{i}$, such that $G_{i}\cong PGL_{2n+1}$ and $G^{\sigma}_{i}\cong P(GL_{n}\times GL_{n+1})$.
  3. 3.
    $\chi^{-1}\chi_{0}$ is trivial on $\Omega_{H}$.
Remark 1.6.
  1. 1.
    Notice that for $\chi=\chi_{0}$ the last condition holds automatically.
  2. 2.
    In this case of $\chi=\chi_{0}$. The opposite direction of Theorem 1.5 also holds. If $St_{\chi_{0}}$ is $H$-distinguished then conditions 1 and 2 hold.
Let $G^{\vee}$ be the dual group of $G$ and let $G^{\vee}_{X}$ be the dual group of $X$ which comes with a map $\iota:G^{\vee}_{X}\times SL_{2}\rightarrow G^{\vee}$ (see [KS17] and [Tak23]).
Conjectuarly (see Conjecture 1.1 in [Tak23]), a representation is $H$ distinguished only if its Langlands parameter factors through $\iota$. We prove a version of this conjecture for the Steinberg representation.
Let $\mathcal{U}$ be the unipotent variety of $G^{\vee}$ and let $\mathcal{U}_{X}$ be the unipotent variety of $G^{\vee}_{X}$.
Theorem 1.7.
Let $F$ be a non archimedean local field. $St_{\chi_{0}}$ is $H$ distinguished if and only if the following equivalent conditions hold:
  1. 1.
    $\iota(\mathcal{U}_{X})$ intersects the open $G^{\vee}$ orbit of $\mathcal{U}$.
  2. 2.
    The Langlands parameter $\phi_{St}$ factors through $G^{\vee}_{X}$.
1.2 Methods of proof
We use the notations and methods of [Sht24].
We consider two hyper graphs related to $X$.
Let $B$ be a Borel subgroup of $G$ and let $\Delta$ be a set of simple roots that corresponds to $B$. For any $\alpha\in\Delta$ let $P_{\alpha}$ be the standard parabolic subgroup that corresponds to $\alpha$.
Definition 1.8.
We define a hyper graph $\Gamma_{F}(X)$ whose set of vertices, $B_{F}(X)$, are the $B$ orbits on $X$ that are defined over $F$. The hyper edges of $\Gamma_{F}(X)$ are labeled by $\Delta$, for any $\alpha\in\Delta$ and $b\in B_{F}(X)$, we have the hyper edge of Borel orbits
$E_{b,\alpha}=\{b^{\prime}\in B_{F}(X)|P_{\alpha}b^{\prime}=P_{\alpha}b\}$
For the case where $F$ is a non archimedean local field, we define a similar graph $\Gamma_{aff,F}(X)$. Let $I$ be an Iwahori subgroup of $G$. Let $\tilde{\Delta}$ be a set of simple reflection in the affine Weyl group $W_{aff}$.
For any $\alpha\in\tilde{\Delta}$ let $I_{\alpha}$ be the parahoric group that contains $I$ and corresponds to $\alpha$.
Definition 1.9.
We define a hyper graph $\Gamma_{aff,F}(X)$ whose set of vertices, $I_{F}(X)$, are the $I$ orbits on $X$. The hyper edges of $\Gamma_{aff,F}(X)$ are labeled by $\tilde{\Delta}$, for any $\alpha\in\tilde{\Delta}$ and $x\in I_{F}(X)$ an $I$ orbit, we have the hyper edge of $I$ orbits $E_{b,\alpha}=\{b^{\prime}\in I_{F}(X)|I_{\alpha}b^{\prime}=I_{\alpha}b\}$.
The group $W_{aff}$ acts on the vertices of the hyper graph $\Gamma_{aff,F}(X)$ (see Theorem 1.1 of [Sht24]). In particular $\Omega\subset W_{aff}$ acts on the vertices of $\Gamma_{aff,F}(X)$. $\Omega$ also acts on the hyper edges of $\Gamma_{aff,F}(X)$.
Denote by $\Gamma^{0}_{aff,F}(X)$ the quotient of $\Gamma_{aff,F}(X)$ by the action of $\Omega$. The vertices of $\Gamma^{0}_{aff,F}(X)$ can be identified with the $H$ orbits of chambers of the extended Bruhat Tits Building $\mathcal{B}$ of $G$.
Definition 1.10.
Let $\Gamma$ be a hyper graph, a function $f:\Gamma\rightarrow\mathbb{C}$ is called harmonic if for every hyper edge $E$, we have $\sum_{e\in E}f(e)=0$. Denote by $\mathcal{H}(\Gamma)$ the space of all harmonic functions on $\Gamma$.
These hyper graphs are related to the problem of determining the distinction of the Steinberg representations by the following proposition:
Proposition 1.11.
  • If $F$ is a finite field then $dim(Hom_{H}(St,1))=dim\mathcal{H}(\Gamma_{F}(X))$.
  • If $F$ is a non archimedean local field then $dim(Hom_{H}(St_{\chi_{0}},1))=dim\mathcal{H}(\Gamma^{0}_{aff,F}(X))$.
In order to show that the Steinberg representation is distinguished it is enough to construct a non zero harmonic functions on the hyper graphs $\Gamma_{F}(X)$ in the finite field case and on $\Gamma_{aff,F}(X)$ in the local field case.
1.3 Structure of the paper
In Section 2, we recall the definition and relevant properties of the Steinberg representation.
In Section 3, we relate the problems about distinction the Steinberg representation to problems about existence of non zero harmonic functions on the hyper graphs $\Gamma_{F}(X)$ and $\Gamma_{aff,F}(X)$.
In Section 4, we prove that $X$ being quasi split is a necessary condition for distinction.
In Section 5, we determine when a non zero harmonic function exists on the hyper graph $\Gamma_{\bar{F}}(X)$. Here $\bar{F}$ is an algebraic closure of $F$.
In Section 6, we discuss harmonic function on $\Gamma_{F}(X)$ for a finite field $F$ and we prove Theorem 1.2.
In Section 7, we discuss harmonic function on $\Gamma_{aff,F}(X)$ we prove Theorem 1.5 for the case $\chi=\chi_{0}$.
In Section 8, we extend the results of Section 7 to a general character and finish the proof of Theorem 1.5.
In Section 9, we relate our results to the dual group $G^{\vee}_{X}$ and prove Theorem 1.7.
1.4 Related works
The problem of distinction of the Steinberg representation in the local field case was studied extensively in the Galois case. Meaning, $E/F$ is a quadratic extension, $G=\mathbf{G}(E)$ and $H=\mathbf{G}(F)$. The work on this case began with the work of Prasad [Pra15]. For specific results see for example [Mat17], [Cho19], [Bro14],[Cou17] and [Bro18]. All the known results are somewhat orthogonal to ours as they deal with non split groups but only in the Galois case.
Interestingly in [Bro18], harmonic functions are also used in order to analyze the distinction of the Steinberg representation, although on different graphs.
Recently a preprint appeared on arXiv dealing with the same problem [WZ24]. In their work they show an upper bound on the dimension of $Hom_{H}(St,\chi)$ and compute it for the case of $G$ being $GL_{n}$ and $H$ being the special orthogonal group. Both our methods rely heavily on the Bruhat Tits building, but otherwise our methods are quite different. They use the description of the Steinberg representation as harmonic cochains on the Bruhat Tits building. We use the description of the Steinberg representation using Type theory and the Hecke Iwahori algebra, as well as the description of $S(X)^{I}$ given in [Sht24].
In some aspects their results are more general than ours as they don’t assume the groups being split.
1.5 Acknowledgement
I would like to thank Raphael Beuzart Plessis for suggesting applying the methods of [Sht24] to the problem of the distinction of the Steinberg representation. I would also like to thank my advisor Eitan Sayag for helpful discussions.
2 The Steinberg representation
The Steinberg is defined both over a finite field field and over a non archimedean local field.
2.1 The Steinberg representation over a finite field
Definition 2.1 (See 1.5 of [Hum87]).
Let $F$ be a finite field and let $\mathcal{B}_{sph}$ be the spherical building of $G$. Let $r_{s}$ be the semi simple rank of $G$. The Steinberg representation, $St$, of $G$ can be defined using the $r-1$ homology group of $\mathcal{B}_{sph}$ with the action of $G$ on it. $St=H_{r-1}(\mathcal{B}_{sph},\mathbb{Q})$.
Proposition 2.2 (See [Hum87]).
$St$ is irreducible.
Proposition 2.3 (See 1.4 of [Hum87]).
Let $F$ be a finite field, let $B$ be a Borel subgroup of $G$ and let ${\cal H}(G,B)$ be the Hecke algebra of $G$. $H(G,B)$ acts on $St^{B}$ by $sgn$. Meaning for every simple reflection $s\in\Delta$, the element $T_{s}=1_{BsB}$ acts on $St^{B}$ by $-1$.
2.2 The Steinberg representation over a non archimedean local field
Definition 2.4.
Let $F$ be a non archimedean local field and let $\mathcal{B}_{aff}$ be the affine building of $G$. The Steinberg representation, $St$, of $G$ is defined to be the top compact cohomology group of $\mathcal{B}_{aff}$ with the action of $G$ on it. $St=H^{top}_{c}(\mathcal{B}_{aff})$
Definition 2.5.
Let $\chi:G\rightarrow\mathbb{C}^{\times}$ be an unramified character. Define $St_{\chi}=St\otimes\chi$.
Proposition 2.6 (See [Bro18]).
$St$ is irreducible.
Let $W_{aff}$ be the affine Weyl group of $G$, let $\Omega$ be the subgroup of $W_{aff}$ of elements of length 00. The group $\Omega$ is embedded in the permutation group of the vertices of the fundamental alcove $\mathcal{C}$. There is a homomorphism $\omega:G\rightarrow\Omega$ which can be defined by comparing the order of the vertices of the alcoves $g\mathcal{C}$ and $\mathcal{C}$. See proposition 2.14 of [Sht24] for a precise definition.
Proposition 2.7 (See subsection 3.3 of [Bro18]).
Let $I$ be an Iwahori subgroup of $G$ and let ${\cal H}(G,I)$ be the affine Hecke algebra of $G$. $St^{I}$ is one dimensional and $H(G,I)$ acts on $St^{I}$ by $sgn$. Meaning for every simple reflection $s\in\tilde{\Delta}$ $T_{s}=1_{IsI}$ acts on $St^{I}$ by $-1$. For $o\in\Omega$, the element $T_{o}=1_{IoI}$ acts on $St^{I}$ by $sgn(o)$, the sign of $o$ as a permutation on the vertices of $\mathcal{C}$.
It follows from the last proposition that $St$ is in the principal block of $G$ representations generated by their $I$ fixed vectors (see [Bor76]).
3 Distinction of the Steinberg representations and harmonic functions
In this Section we will relate the problem about determining the distinction of the Steinberg representation to the problem of existence of a non zero harmonic function on the hyper graphs $\Gamma_{F}(X)$ and $\Gamma_{aff,F}(X)$ defined in definitions 1.8 and 1.9 respectively.
We begin with the finite field case.
Proposition 3.1.
If $F$ is a finite field then $dim(Hom_{H}(St,1))=dim\mathcal{H}(\Gamma_{F}(X))$.
Proof.
By Frobenius reciprocity we have $Hom_{H}(St,1)=Hom_{G}(St,ind^{G}_{H}(1))=Hom_{G}(St,\mathbb{C}[X])$, Here $\mathbb{C}[X]$ is the representation of functions on $X$.
Let $B$ be a Borel subgroup of $G$. $St$ is irreducible and has a $B$ fixed vector. We can write $\mathbb{C}[X]$ as a direct sum of its sub representation generated by its $B$ fixed vector and some complement $\mathbb{C}[X]=G\mathbb{C}[X]^{B}\oplus\pi$. Any map from $St$ to $\mathbb{C}[X]$ will land in $G\mathbb{C}[X]^{B}$, so $Hom_{G}(St,\mathbb{C}[X])=Hom_{G}(St,G\mathbb{C}[X]^{B})$.
Both the Steinberg representation and $G\mathbb{C}[X]^{B}$ are self dual representations of $G$, therefore $Hom_{G}(St,G\mathbb{C}[X]^{B})=Hom_{G}(G\mathbb{C}[X]^{B},St)$.
$\mathbb{C}[X]^{B}$ and $St^{B}$ are representation of the Hecke algebra $H(G,B)$ of $B$ bi-onvariant functions on $G$ and we have $Hom_{G}(G\mathbb{C}[X]^{B},St)=Hom_{{\cal H}(G,B)}(\mathbb{C}[X]^{B},St^{B})$.
We will show that $Hom_{{\cal H}(G,B)}(\mathbb{C}[X]^{B},St^{B})=\mathcal{H}(\Gamma_{F}(X))$.
The module $St^{B}$ is one dimensional. An element of $Hom_{{\cal H}(G,B)}(\mathbb{C}[X]^{B},St^{B})$ is a function on the $B$ orbits on $X$.
Let $f\in Hom_{{\cal H}(G,B)}(\mathbb{C}[X]^{B},St^{B})$, we know that for every $\alpha\in\Delta$, $T_{\alpha}f=-f$. Thus for every Borel orbit $b\in B\backslash X$ we have $((T_{s}+1)f)(b)=0$, this is exactly the condition $\sum_{b^{\prime}\in E_{b,\alpha}}f(b^{\prime})=0$ which means that $f$ is a harmonic function on $\Gamma_{F}(X)$. ∎
Now we move to the local field case. The proof is almost the same.
Proposition 3.2.
If $F$ is a non archimedean local field, $dim(Hom_{H}(St_{\chi_{0}},1))=dim\mathcal{H}(\Gamma^{0}_{aff,F}(X))$.
Proof.
By Frobenius reciprocity we have
$Hom_{H}(St_{\chi_{0}},1)=Hom_{G}(St_{\chi_{0}},ind^{G}_{H}(1))=Hom_{G}(St_{\chi_{0}},C^{\infty}(X))$
Here $C^{\infty}(X)$ is the representation of locally constant functions on $X$.
The representation $St_{\chi_{0}}$ is self dual, therefore $Hom_{G}(St_{\chi_{0}},C^{\infty}(X))=Hom_{G}(S(X),St_{\chi_{0}})$. Here $S(X)$ is the representation of locally constant compactly supported functions on $X$.
Let $I$ be an Iwahori subgroup of $G$. $St_{\chi_{0}}$ is irreducible and is generated by its $I$ fixed vectors. The subgroup $I$ splits the category of smooth representations of $G$ (see [Ber84]), and the category of representations generated by their $I$ fixed vectors is equivalent to the category of $\mathcal{H}(G,I)$ modules (see [Mat77]). Therefore $Hom_{G}(S(X),St_{\chi_{0}})=Hom_{{\cal H}(G,I)}(S(X)^{I},St_{\chi_{0}}^{I})$.
We will show that $Hom_{{\cal H}(G,I)}(S(X)^{I},St_{\chi_{0}}^{I})=\mathcal{H}(\Gamma^{0}_{aff,F}(X))$. The ${\cal H}(G,I)$ module $St_{\chi_{0}}^{I}$ is one dimensional so an element of $Hom_{{\cal H}(G,I)}(S(X)^{I},St_{\chi_{0}}^{I})$ is a function on the $I$ orbits on $X$.
Let $f\in Hom_{{\cal H}(G,I)}(S(X)^{I},St_{\chi_{0}}^{I})$, we know that for every $s\in\tilde{\Delta}$, $T_{s}f=-f$. Thus for every $I$ orbit $b\in I\backslash X$ we have $((T_{s}+1)f)(b)=0$, this is exactly the condition $\sum_{b^{\prime}\in E_{b,s}}f(b^{\prime})=0$ (see Section 4 of [Sht24]) which means that $f$ is a harmonic function on $\Gamma^{0}_{aff,F}(X)$.
The elements $T_{s}$ for $s\in\tilde{\Delta}$ might not generate ${\cal H}(G,I)$, we also have to consider the action of the elements $T_{o}$ for $o\in\Omega$. We have $(T_{o}f)(b)=\chi^{2}_{0}(o)f(b)=f(b)$. On the other hand, $(T_{o}f)(b)=f(o^{-1}\times b)$ (see proposition 4.1 in [Sht24]). Therefore, $f$ is constant on $\Omega$ orbits.
Any harmonic function on $\Gamma^{0}_{aff,F}(X)$ can be extended in a unique way to a harmonic function on $\Gamma_{aff,F}(X)$ which is constant on $\Omega$ orbits. Thus we are done.
In the previous proposition we related the question about the distinction of $St_{\chi_{0}}$ to a question about the existence of non zero harmonic functions. Next we want to relate the question about the distinction of $St_{\chi}$ for a general unramified character $\chi:G\rightarrow\mathbb{C}^{\times}$ to a question about the existence of non zero harmonic functions.
The vertices of $\Gamma_{aff,F}(X)$ can be identified with the set of $H$ orbits on $\Omega$ colored chambers in the extended Bruhat Tits building of $G$, which we denote by $\mathcal{B}$ (see Section 3 of [Sht24]). The vertices of $\Gamma^{0}_{aff,F}(X)$ can be identified with the set of $H$ orbits on the chambers in $\mathcal{B}$ as we divide by the action of $\Omega$.
We introduce one more hyper graph. Let $\Gamma_{aff,F}^{1}(X)$ be the sub hyper graph of $\Gamma_{aff,F}(X)$ whose vertices correspond to $H$ orbits of chambers colored by the unit $1\in\Omega$. The vertices of $\Gamma_{aff,F}^{1}(X)$ can be identified with $H\cap Ker(\omega)$ orbits on the chambers of $\mathcal{B}$. Denote $H^{1}=H\cap Ker(\omega)$.
Proposition 3.3.
$St_{\chi}$ is $H$ distinguished if and only if there is a non zero $\phi\in\mathcal{H}(\Gamma^{1}_{aff,F}(X))$ such that for any chamber $\mathcal{C}$ and $h\in H$, $\phi(H^{1}h\mathcal{C})=\phi(H^{1}\mathcal{C})\chi^{-1}\chi_{0}(\omega(h^{-1}))$.
Proof.
Like in the proof of proposition 3.2 we have $Hom_{H}(St_{\chi},1)=Hom_{G}(S(X),St_{\chi^{-1}})$ which can described as a subspace of the space of harmonic functions $\phi$ on $\Gamma_{aff,F}(X)$. The extra condition on $\phi$ to belong to this subspace is the following. Let $o\in\Omega$ and $\mathcal{C}\subset\mathcal{B}$ be a chamber. Denote by $x_{\mathcal{C},o}$ the $I$ orbit on $X$ the correspondence to $\mathcal{C}$ with the color $o$. The condition is $\phi(x_{\mathcal{C},o})=\chi_{0}(o)\chi^{-1}(o)\phi(x_{\mathcal{C},1})$.
For the first direction, assume that there exists a non zero harmonic function $\phi$ on $H$ orbits of $\Omega$ colored chambers satisfying the extra condition above. Restricting it to chambers with the color $1\in\Omega$ and identifying this set with the set of $H\cap Ker(\omega)$ orbits on the chambers of $\mathcal{B}$ will give us a harmonic function on $\Gamma^{1}_{aff,F}(X)$.
Clearly, if $h\in H$ and $\mathcal{C}$ a chamber then
$\phi(H^{1}h\mathcal{C})=\phi(x_{h\mathcal{C},1})=\phi(x_{\mathcal{C},\omega(h^{-1})})=\chi_{0}\chi^{-1}(\omega(h^{-1}))\phi(x_{\mathcal{C},1})=\chi_{0}\chi^{-1}(\omega(h^{-1}))\phi(H^{1}\mathcal{C})$
For the second direction, assume that there exists a non zero harmonic function $\phi$ on $\Gamma^{1}_{aff,F}(X)$ such that chamber for any chamber $\mathcal{C}$ and $h\in H$, $\phi(H^{1}h\mathcal{C})=\phi(H^{1}\mathcal{C})\chi^{-1}\chi_{0}(\omega(h^{-1}))$. We define a function $\phi^{\prime}$ on $\Gamma_{aff,F}(X)$ by assigning to a $H$ orbit of a chamber $\mathcal{C}$ with the color $o$ the value $\phi^{\prime}(\mathcal{C},o)=\phi(H^{1}\mathcal{C})\chi^{-1}\chi_{0}(o)$. We only need to check that this function is $H$ invariant. Let $h\in H$, $o\in\Omega$ and $\mathcal{C}$ a chamber, $\phi^{\prime}(h\mathcal{C},\omega(h)o)=\phi(H^{1}h\mathcal{C})\chi^{-1}\chi_{0}(\omega(h)o)=\phi(H^{1}\mathcal{C})\chi^{-1}\chi_{0}(o)=\phi^{\prime}(\mathcal{C},o)$. ∎
4 $X$ being quasi split is a necessary condition for distinction
In this section we will prove that $X$ being quasi split is a necessary condition for distinction in the non archimedean local field case. With slight modifications, our proof can be applied also to the finite field case.
Remark 4.1.
The results actually follows from known results. It is known that $St$ is generic, i.e. has a Whittaker model (see for example [BBBG24]). By [Pra19] a generic representation can be distinguished only if $X$ is quasi split.
We provide a more direct proof.
Let $F$ be a non archimedean local field and let $\mathcal{B}$ be the affine building of $G$.
Recall that $\sigma$ acts on $\mathcal{B}$ and that an apartment, $\mathcal{A}$ , of $\mathcal{B}$ is called $\sigma$ stable if $\sigma(\mathcal{A})=\mathcal{A}$.
We introduce a partial order on the $\sigma$ stable apartments.
Definition 4.2.
Let $\mathcal{A}$ be a $\sigma$ stable apartment, and let $T\subset G$ be a maximal split torus that corresponds to $\mathcal{A}$. Define $rank(\mathcal{A})=rank(T)-rank(T\cap H)$.
Definition 4.3.
We define a partial order $<$ to be the minimal partial order such that for any two $\sigma$ stable apartments $\mathcal{A}_{1},\mathcal{A}_{2}$ with $rank(\mathcal{A}_{1})<rank(\mathcal{A}_{2})$ and $\mathcal{A}_{1}\cap\mathcal{A}_{2}$ is of codimension 1 inside $\mathcal{A}_{1}$ (equivalently $\mathcal{A}_{2}$) we have $\mathcal{A}_{1}<\mathcal{A}_{2}$.
This partial order descends to a partial order on the $H$ orbits of $\sigma$ stable apartments.
Proposition 4.4.
Let $\mathcal{A}_{1},\mathcal{A}_{2}$ be two $\sigma$ stable apartments such that $H\mathcal{A}_{1}\neq H\mathcal{A}_{2}$ and such that their intersection is of codimesnion 1 inside them. Then $|rank(\mathcal{A}_{1})-rank(\mathcal{A}_{2})|\leq 1$. Furthermore, $\mathcal{A}_{1}<\mathcal{A}_{2}$ if and only if the action of $\sigma$ on $\mathcal{A}_{1}$ preserves the two half planes defined by $\mathcal{A}_{1}\cap\mathcal{A}_{2}$.
Proof.
Let $\mathcal{A}$ be a $\sigma$ stable apartment, $rank(\mathcal{A})$ is equal to the co-dimension of the subgroup of translations on the affine space $\mathcal{A}$ which commute with the action of $\sigma$ on $\mathcal{A}$ (see proposition 5.7 in [Sht24]).
Now, let $\mathcal{A}_{1},\mathcal{A}_{2}$ be two $\sigma$ stable apartments whose intersection is of codimesnion 1. Denote by $d$ the dimension of the group of transaltion on $\mathcal{A}_{1}\cap\mathcal{A}_{2}$ that commute with $\sigma$.
Denote $n=dim\mathcal{A}_{1}=dim\mathcal{A}_{2}$. Then, for $i\in\{1,2\}$, $rank(\mathcal{A}_{i})$ is either $n-d$ or $n-d-1$. If The translation in $\mathcal{A}_{i}$ which is perpendicular to $\mathcal{A}_{1}\cap\mathcal{A}_{2}$ commutes with $\sigma$ then $rank(\mathcal{A}_{i})=n-d-1$, otherwise $rank(\mathcal{A}_{i})=n-d$.
Let $\mathcal{C}$ be a chamber in $\mathcal{A}_{i}$ that contains a face of codimesnion 1 that spans $\mathcal{A}_{1}\cap\mathcal{A}_{2}$ (as an affine hyper plane). The translation perpendicular to $\mathcal{A}_{1}\cap\mathcal{A}_{2}$ commutes with $\sigma$ if and only if $\sigma(\mathcal{C})$ and $\mathcal{C}$ are on the same side of $\mathcal{A}_{1}\cap\mathcal{A}_{2}$. By Remark 7.9 of [Sht24], this occurs if and only if $\mathcal{C}$ has minimal $l_{\sigma}$ length among the chambers that contain the face $f=\mathcal{C}\cap\mathcal{A}_{1}\cap\mathcal{A}_{2}$.
We know that there is at most one $H$ orbit of $\sigma$ stable apartments that contain a chamber $\mathcal{C}$ that satisfies the above property (see Section 7 of [Sht24]).
Assume that $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ aren’t in the same $H$ orbit, then there is a chamber $\mathcal{C}\subset\mathcal{A}_{1}$ that satisfies the above property if and only if $\mathcal{A}_{1}<\mathcal{A}_{2}$. This finishes the proof. ∎
Definition 4.5.
We say that $X$ is quasi split if there is a Borel subgroup, $B$, of $G$ such that $B\cap\sigma(B)$ is a torus.
We will use an equivalent condition which we now state.
Definition 4.6.
A torus $T\subset G$ is called $\sigma$ split if it is $\sigma$ stable and $\sigma$ acts on it by inverse.
Proposition 4.7 ([Vus74]).
$X$ is quasi split if and only if there is a maximal $\sigma$ split torus $T^{-}$ such that $Z_{G}(T^{-})$ is a maximal torus of $G$.
By proposition 3.2 the fact that $X$ being quasi split is a necessary condition for the distinction of $St_{\chi_{0}}$ is equivalent to the next theorem, in fact by the proof of 3.2 the next theorem implies that $X$ being quasi split is a necessary condition for the distinction of $St_{\chi}$ for any unramified character $\chi$.
Theorem 4.8.
If $X$ isn’t quasi split then $\mathcal{H}(\Gamma^{0}_{aff,F}(X))=0$.
In order to prove the above theorem we need the following proposition:
Proposition 4.9.
Assume that $X$ isn’t quasi split. For every $\sigma$ stable torus $T$ there is a root $\alpha:T\rightarrow F^{\times}$ such that both $u_{\alpha}(x)$ and $u_{-\alpha}(x)$ are in $H$ for every $x\in F$. Here is the homomorphism $u_{\alpha}:\mathbb{G}_{a}\rightarrow\mathbf{G}$.
Proof.
We prove this by induction on the partial order defined by definition 4.3. We begin with the the maximal tori, which are split tori $T$ such that $T^{-}$ is a maximal $\sigma$ split torus.
By our assumption $X$ isn’t quasi split, so $M=Z_{G}(T^{-})$ isn’t a torus. $M$ is a Levi subgroup of $G$ (see for example [Spr09]). Let $\alpha^{\vee}:F^{\times}\rightarrow T$ be a coroot whose unipotent subgroup $U_{\alpha}$ is contained in $M$. Notice that $\sigma(\alpha^{\vee})=\alpha^{\vee}$ as for $t\in T^{-}$ and $x\in F$, $u_{\alpha}(x)=t^{-1}u_{\alpha}(x)t=u_{\alpha}(\alpha(t)x)$ so $\alpha(t)=1$ for $t\in T^{-}$. Thus $\sigma(\alpha)$ and $\alpha$ are equal on $T^{-}$ and they are clearly equal on the $\sigma$ fixed part.
The same argument also works for $-\alpha$.
We claim that for every $x\in F$ $u_{\alpha}(x)\in H$. We know that either $\sigma(u_{\alpha}(x))=u_{\alpha}(x)$ or $\sigma(u_{\alpha}(x))=u_{\alpha}(-x)$, as we showed that $\alpha$ is fixed by $\sigma$. We assume by contradiction that for every $x\in F$, $\sigma(u_{\alpha}(x))=u_{\alpha}(-x)$. It follows that also for every $x\in F$, $\sigma(u_{-\alpha}(x))=u_{-\alpha}(-x)$.
Let $y=u_{\alpha}(\frac{1}{2})u_{-\alpha}(-1)$. We will show that the torus $y^{-1}Ty$ contains a $\sigma$ split torus larger than $T^{-}$ and this will be a contradiction. We have $y\in M$ so clearly $y^{-1}Ty$ contains $T^{-}$. It is enough to show that for every $x\in F^{\times}$, $\sigma$ acts as inverse on $y^{-1}\alpha^{\vee}(x)y$.
We need to show that
$u_{-\alpha}(-1)u_{\alpha}(\frac{1}{2})\alpha^{\vee}(x^{-1})u_{\alpha}(-\frac{1}{2})u_{-\alpha}(1)=u_{-\alpha}(1)u_{\alpha}(-\frac{1}{2})\alpha^{\vee}(x)u_{\alpha}(\frac{1}{2})u_{-\alpha}(-1)$
This is equivalent to showing
$u_{\alpha}(\frac{1}{2})u_{-\alpha}(-2)u_{\alpha}(\frac{1}{2})\alpha^{\vee}(x^{-1})u_{\alpha}(-\frac{1}{2})u_{-\alpha}(2)u_{\alpha}(-\frac{1}{2})=\alpha^{\vee}(x)$
Using $u_{\alpha}(\frac{1}{2})u_{-\alpha}(-2)u_{\alpha}(\frac{1}{2})=n_{\alpha}\alpha^{\vee}(\frac{1}{2})$ where $n_{\alpha}=u_{\alpha}(1)u_{-\alpha}(-1)u_{\alpha}(1)$ (see 8.1.4 in [Spr09]) we are left with the known result (again, see 8.1.4 in [Spr09]):
$n_{\alpha}\alpha^{\vee}(x^{-1})n_{\alpha}^{-1}=\alpha^{\vee}(x)$
Thus we proved the proposition for tori which are maximal with respect to our partial order.
Now it is to prove that if we have two apartments $\mathcal{A}_{1}<\mathcal{A}_{2}$, with $\mathcal{A}_{1}\cap\mathcal{A}_{2}$ of codimension 1 inside $\mathcal{A}_{2}$. Let $T_{1}$, $T_{2}$ be the corresponding maximal split tori to $\mathcal{A}_{1}$,$\mathcal{A}_{2}$ respectively. Assume that we know the proposition for $T_{2}$, then it is true also for $T_{1}$.
Let $\alpha$ be a root of $T_{2}$ such that for every $x\in F$, $u_{\alpha}(x),u_{-\alpha}(x)\in H$. Let $\beta$ be a root of $T_{2}$ which is constant on $\mathcal{A}_{1}\cap\mathcal{A}_{2}$. There is an element $g$ in the group generated by $\{u_{\beta}(y),u_{-\beta}(y)|y\in F\}$ such that $g\mathcal{A}_{2}=\mathcal{A}_{1}$. It is enough to show that $g$ commutes with $u_{\pm\alpha}(x)$, as the conjugation of $\alpha$ by $g$ is a root of $T_{1}$, and if $g$ commutes with $u_{\pm\alpha}(x)$ then $g^{-1}u_{\pm\alpha}(x)g=u_{\pm\alpha}(x)\in H$. It is enough to show that $\alpha,\beta$ are perpendicular.
By proposition 4.4, $\sigma$ acts by $-1$ on $\beta$. It acts as 1 on $\alpha$, therefore $<\alpha,\beta>=<\sigma(\alpha),\sigma(\beta)>=-<\alpha,\beta>$ and we are done.
Definition 4.10.
Let $\mathcal{A}$ be some $\sigma$ stable apartment that corresponds to a maximal split torus $T$. Let $\alpha$ be the root on $T$ given by the previous proposition. We call a chamber $\mathcal{C}$ in $\mathcal{A}$ $\alpha$ adjacent if it contains a face of codimension 1 which is part of a wall defined by an affine translation of $\alpha$. We call a chamber $H$-wall adjacent if its $\alpha$ adjacent for some $\alpha$ that satisfies the property of the previous definition.
Notice that if $h\in H$ acts on $\mathcal{A}$ then it acts on the set of $H$-wall adjacent chambers. Also $\sigma$ acts on the set of $H$-wall adjacent chambers.
Now we can prove Theorem 4.8.
Proof of Theorem 4.8.
Let $\phi\in\mathcal{H}(\Gamma^{0}_{aff,F}(X))$, we will show that it is zero. Recall that $I$ orbits on $X$ can be identified with $H$ orbits of $\Omega$ colored chambers in $\mathcal{B}$ (see Section 3 in [Sht24]). We are concerned only with the connected component of $1$, so we may look only at chambers with the color $1\in\Omega$. We think about $\phi$ as a function on the chambers in $\mathcal{B}$. We will prove that $\phi$ is zero on all chambers.
Clearly $\phi$ is zero on all $\alpha$ adjacent chambers as the edge labeled by $\alpha$ which contains an $\alpha$ adjacent chamber has size 1.
All the walls defined by $\alpha$ are perpendicular to the fixed points of $\sigma$ on $\mathcal{A}$.
For every chamber $\mathcal{C}$ in $\mathcal{A}$ we choose a gallery $(\mathcal{C}_{i})^{n}_{i=1}$ in $\mathcal{A}$, with $\mathcal{C}_{1}$ being $H$-wall adjacent and $\mathcal{C}_{n}=\mathcal{C}$ and $n$ being minimal.
We will prove that $\phi(\mathcal{C}_{i})=0$ for every $i$. We have already discussed this for $i=1$. It is enough to show that $\phi(\mathcal{C}_{i})+\phi(\mathcal{C}_{i+1})=0$.
Notice that for every $i$ $H\mathcal{C}_{i}\neq H\mathcal{C}_{i+1}$ otherwise we could have find a shorter gallery.
If the hyper edge that contains the $H$ orbits of $\mathcal{C}_{i}$ and $\mathcal{C}_{i+1}$ is of size 2, then $\phi(\mathcal{C}_{i})+\phi(\mathcal{C}_{i+1})=0$. Assume it is of size larger than 2, then there is another $\sigma$ stable apartment $\mathcal{A}^{\prime}$ that contains $\mathcal{C}_{i}\cap\mathcal{C}_{i+1}$, we claim that $\mathcal{A}_{1}<\mathcal{A}_{2}$. By propitiation 4.4 it is enough to show that $\sigma$ preserves the two half planes defined by $\mathcal{C}_{i}\cap\mathcal{C}_{i+1}$, otherwise we can find a $H$-wall adjacent chamber which is closer to $\mathcal{C}_{i+1}$ than $\mathcal{C}_{1}$, a contradiction.
Now we use induction on the rank of $\mathcal{A}$, by the previous argument for $\mathcal{A}$ of maximal rank we are done.
For $\mathcal{A}$ not necessarily of maximal rank, we know by induction that $\phi$ is zero on all chambers in $\mathcal{A}^{\prime}$, thus its sum over all chamber in the hyper edge of $\mathcal{C}_{i}$ and $\mathcal{C}_{i+1}$ is 0, and it 0 on all other chambers. Therefore $\phi(\mathcal{C}_{i})+\phi(\mathcal{C}_{i+1})=0$. ∎
5 Harmonic functions in the algebraically closed field case
In the previous section we proved a necessary condition for the existence of a non zero harmonic function on $\Gamma_{aff,F}(X)$. In this section we begin working on the other direction, which requires constructing harmonic functions on our hyper graphs. Let $\bar{F}$ be the algebraic closure of $F$. In this section we will construct a non zero harmonic function on $\Gamma_{\bar{F}}(X)$ under the assumption that $X$ is quasi split. In the following sections we will upgrade this construction to $\Gamma_{aff,F}(X)$ and $\Gamma_{F}(X)$.
Definition 5.1.
Let $\Gamma$ be a hyper graph whose hyper edges are labeled by a finite set $L$, let $L^{\prime}\subset L$. We call a vertex $v$ of $\Gamma$, $L^{\prime}$ full if for every label $\alpha\in L^{\prime}$ the hyper edge with label $\alpha$ that contains $v$ is of size at least two. In the case $L^{\prime}=L$ we call an $L$ full vertex just full.
The following is a result of Prasad [Pra19].
Proposition 5.2.
(Proposition 9 of [Pra19]) $X$ is quasi split if and only if there is a closed Borel orbit which is full in $\Gamma_{\bar{F}}(X)$.
Remark 5.3.
A natural question to ask is if the previous result holds also for spherical varieties which are not symmetric. This will be a first step in extending our results to the general case of general spherical varieties. In the general spherical variety case the notion of being quasi split is replaced with the notion of being tempered, which means that the stabilizer of the open Borel orbit on $X$ is a Borel subgroup of $G$.
We will prove the following result:
Proposition 5.4.
The number of full closed Borel orbits in $\Gamma_{\bar{F}}(X)$ is equal to $dim\mathcal{H}(\Gamma_{\bar{F}}(X))$.
Using propositions 5.2 and 5.4 it follows that $X$ is qausi split if and only if there is a non zero harmonic function on $\Gamma_{\bar{F}}(X)$.
Denote by $S$ the set of full closed Borel orbits in $\Gamma_{\bar{F}}(X)$. The main component in the proof of proposition 5.2 is the following:
Proposition 5.5.
There exists a function $s:B\backslash X\rightarrow\mathbb{Z}[S]$ such that:
  1. 1.
    For any $x\in S$ we have $s(x)=x$.
  2. 2.
    For any $E$, a hyper edge of $\Gamma_{\bar{F}}(X)$, $\sum_{v\in E}s(v)=0$.
Proof.
We construct $s$ by induction on the Bruhat order on $B\backslash X$. It is defined on the full closed orbits by the first condition and we define it to be zero on the non full closed orbits.
Let $v$ be a Borel orbit and assume that $s$ was already defined on all orbits which are less than $v$ in the Bruhat order. Let $\alpha,\beta\in\Delta$ be two roots such that the hyper edges containing $v$ defined by them contain only orbits which are equal or less than $v$ (in the bruhat order). There are four options for each of them, they can be roots of type $G,T,U$ or $N$ (see [Kno95]).
Let $\Gamma_{v,\{\alpha,\beta\}}$ be the sub hyper graph containing hyper edges labeled either $\alpha$ or $\beta$ and vertices which are connected to $v$ by hyper edges labeled $\alpha$ or $\beta$. This is a hyper graph of a spherical variety of the rank 2 Levi subgroup that corresponds to the roots $\alpha,\beta$ (see Lemma 3 of [Bri01]). It is enough to show the proposition for rank 2 groups because it will prove that defining $s$ using condition 2 and the Borel orbit in the hyper edge labeled by $\alpha$ will give the same result as defining $s$ using condition 2 and the Borel orbits in the hyper edge labeled by $\beta$.
Thus it is enough to check the proposition for all spherical varieties of group of rank 2. In order to minimize the case by case examination we will prove that it is enough to check the proposition for symmetric spaces. In proposition 7.2 we will prove an affine version of this reduction (see remark 7.3). The same method can be used in the algebraically closed field setting, switching the affine building for the spherical one.
We will further reduce to adjoint groups.
Let $\mathbf{Z}$ be the center of $\mathbf{G}$ and let $\mathbf{G}^{adj}=\mathbf{G}/\mathbf{Z}$ be the adjoint form of $G$. Let $\pi_{adj}:\mathbf{G}\rightarrow\mathbf{G}^{adj}$ be the natural projection on the level of algebraic groups. The involution $\sigma$ acts on $\mathbf{G}^{adj}$. The fixed points of $\sigma$ on $\mathbf{G}^{adj}$ might be larger than $\pi_{adj}(\mathbf{H})$.
We claim that $\pi_{adj}(\mathbf{H})$ must contain the connected component of $1$ in $(\mathbf{G}^{adj})^{\sigma}$. We show this on the level of $\bar{F}$ points. We write $G^{adj}=\mathbf{G}^{adj}(\bar{F})$ and $Z=\mathbf{Z}(\bar{F})$
Any $g\in G^{adj}$ can be lifted to an element of $g^{\prime}\in G^{\prime}$, the derived group of $G$. The group $G^{\prime}\cap Z$ is finite and if $g\in(G^{adj})^{\sigma}$ then $g^{\prime}\sigma(g)^{-1}\in G^{\prime}\cap Z$ which is finite, therefore on the connected component of $1$ it is constantly $1$ and $g^{\prime}\in H$.
Denote by $H^{0}$ the connected component of $1\in H$. The group $N_{G}(H^{0})/H^{0}$ acts on $\Gamma_{\bar{F}}(G/H^{0})$ and it contains $H/H^{0}$. $\Gamma_{\bar{F}}(G/H)$ is the quotient of $\Gamma_{\bar{F}}(G/H^{0})$ with respect to the action of $H/H^{0}$.
Thus we have to verify the proposition for symmetric subgroups and their connected components of $PGL_{2}\times PGL_{2},PGL_{3},PSp_{4}$ and $G_{2}$.
The case of $PGL_{2}\times PGL_{2}$: The symmetric subgroups of $PGL_{2}\times PGL_{2}$ are either conjugate to the diagonal $PGL_{2}$ or are a product of two symmetric subgroups of $PGL_{2}$. In the former case the hyper graph $\Gamma_{\bar{F}}(X)$ has two vertices which are connected by two edges labeled $\alpha$ and $\beta$. In the latter case $\Gamma_{\bar{F}}(X)$ is a product of two hyper graphs of $PGL_{2}$. In any case it is easy to see that the proposition holds.
The case of $PGL_{3}$: Up to isomorphism the involutions on $PGL_{3}$ are $x\rightarrow(x^{t})^{-1}$ and $Ad(t)$ where $t$ is a diagonal matrix with two $-1$ and on $1$ on the diagonal. The connected components of their fixed points are $PO(3)$ and $PGL_{2}$. The hyper graphs that correspond to these groups can be see in figure 1. In both cases there is a unique full closed orbits and $s$ is well defined. This is true also when considering all possible quotient of these hyper graphs.
[…]
Figure 1: On the left the hyper graph $\Gamma_{\bar{F}}(PGL_{3}/PO_{3})$ and on the right the hyper graph $\Gamma_{\bar{F}}(PGL_{3}/PGL_{2})$. $\alpha,\beta$ are the two simple roots of $PGL_{3}$
The case of $PSp_{4}$: In this case, up to isomorphism there is a unique involution which is an inner involution. The connected component of the group of its fixed points is isomorphic to $PGL_{2}$ and its $\Gamma_{\bar{F}}(X)$ hyper graph is given in figure 2, see also example 3 in [Bri01]. It has a unique quotient, which also appears in example 3 in [Bri01]. For both this hyper graph and its quotient, the function $s$ is well defined.
[…]
Figure 2: The hyper graph $\Gamma_{\bar{F}}(PSp_{4}/PGL_{2})$. $\alpha,\beta$ are the two simple roots of $PSp_{4}$.
The case of $G_{2}$: In this case, up to isomorphism there is a unique involution which is an inner involution. The group of its fixed points is isomorphic to $PGL_{2}\times SL_{2}$. Its hyper graph is given in figure 3. For this hyper graph the function $s$ is well defined.
[…]
Figure 3: The hyper graph $\Gamma_{\bar{F}}(G_{2}/PGL_{2}\times SL_{2})$. $\alpha,\beta$ are the two simple roots of $G_{2}$.
Now we can prove proposition 5.4.
Proof of proposition 5.4.
By composing the map $s$ from the previous proposition with any function $f:\mathbb{Z}[S]\rightarrow\mathbb{C}$ we obtain a harmonic function on $\Gamma_{\bar{F}}(X)$. Thus the dimension of harmonic functions is at least the number of closed full orbits. If the dimension of harmonic functions is larger than the number of full closed orbits then there is a harmonic function which is zero on all full closed orbits. Such a function is zero on all the closed orbits. We claim that such a function must be zero on the entire hyper graph. This follows from induction on the dimension of the orbits and from the fact that each hyper edge can be described using a spherical variety of $PGL_{2}$ (see [Kno95]).
Any spherical variety of $PGL_{2}$ has a unique non closed orbit (again see [Kno95]) and thus a harmonic function which is zero on all closed orbits must be zero.
6 Harmonic functions in the finite field case
In this Section we will construct a non zero harmonic function on $\Gamma_{F}(X)$ in the case of $F$ being a finite field, $X$ being quasi split and $G,H$ are both split over $F$. Our main tools are proposition 5.4 of the previous section and the following result of [Hun21].
Theorem 6.1 ([Hun21]).
Assume that both $G$ and $H$ are split over $F$, the natural map $\pi:\Gamma_{F}(X)\rightarrow\Gamma_{\bar{F}}(X)$ is surjective.
Proposition 6.2.
Let $v$ be a dense $B$ orbit of $X$ over $F$. Consider $\Gamma^{v}_{F}(X)$, the sub hyper graph of $\Gamma_{F}(X)$ which consists of vertices below $v$. The natural map $\pi:\Gamma^{v}_{F}(X)\rightarrow\Gamma_{\bar{F}}(X)$ is an isomorphism.
Proof.
We prove it by induction on the Bruhat order. On the level of the dense orbit it is clear. Now, assume we have some orbit $x\in\Gamma^{v}_{F}(X)$ and we have some root $\alpha$ and $y\in\Gamma_{\bar{F}}(X)$ such that $y$ and $\pi(x)$ are in a hyper edge labeled by $\alpha$. It is enough to show that there is a unique orbit $y^{\prime}\in\Gamma^{v}_{F}(X)$ such that $\pi(y^{\prime})=y$ and $y^{\prime},x$ are in a hyper edge labeled by $\alpha$.
The hyper edge labeled by $\alpha$ correspondence to Borel orbits of a spherical verity of a subgroup of $G$ of semi-simple rank 1 (this is the Levi of $P_{\alpha}$, a parabolic that correspondence to $\alpha$). This subgroup can be constructed as centralizer of some torus, and then corollary 5.17 of [Hun21] can be applied to find $y^{\prime}$.
When we restrict $\pi$ to a single hyper edge it is injective on the set of orbits of minimal dimension in the hyper edge, this is because the hyper edge correspondence to Borel orbits on a spherical variety of $PGL_{2}$. And for any of such variety passing to algebraic closer is injective on closed orbits (see [Kno95] for the description of these varieties and their Borel orbits).
What is left to show, is that $\pi$ can’t glue two different hyper edges. This follows from the rationality of Knop’s action (see proposition 3.6.1 in [Sak08]). ∎
Definition 6.3.
Define a partial order on $\Gamma_{F}(X)$ by taking the minimal partial order such that for every two Borel orbits $x_{1},x_{2}$ in the same hyper edge, $x_{1}<x_{2}$ if $\overline{x_{1}}\subsetneq\overline{x_{2}}$.
Definition 6.4.
Let $x$ be a vertex of $\Gamma_{F}(X)$, denote by $m(x)$ the number of dense orbits above it in the partial order defined above.
Let $\phi:\Gamma_{\bar{F}}(X)\rightarrow\mathbb{C}$, define $\pi^{!}(\phi):\Gamma_{F}(X)\rightarrow\mathbb{C}$ by $\pi^{!}(\phi)(x)=m(x)\phi(\pi(x))$.
Proposition 6.5.
For $\phi\in\mathcal{H}(\Gamma_{\bar{F}}(X))$ we have $\pi^{!}(\phi)\in\mathcal{H}(\Gamma_{F}(X))$.
Proof.
Let $B$ be a Borel subgroup of $G$. Let $\mathcal{O}$ be a $B$ orbit on $X$ over $F$. Let $\alpha\in\Delta$ be some root and let $P_{\alpha}$ be the corresponding standard parabolic that contains $B$. Write $P_{\alpha}\mathcal{O}=\mathcal{O}_{1}\cup\mathcal{O}_{1}...\cup\mathcal{O}_{k}$ as a union of $B$ orbits. We want to show that $\sum^{k}_{i=1}m(\mathcal{O}_{i})\phi(\pi(\mathcal{O}_{i}))=0$. It is enough to show that for every orbit $\mathcal{O}_{l}$, $\sum_{\pi(\mathcal{O}_{i})=\pi(\mathcal{O}_{l})}m(\mathcal{O}_{i})$ doesn’t depend on $\mathcal{O}_{l}$. From calculating what happens in the case of $G=PGL_{2}$, we know that if we restrict $\pi$ to $\mathcal{O}_{1},...,\mathcal{O}_{k}$ then it is injective on the subset of orbits closed in $P_{\alpha}\mathcal{O}$ and it maps all orbits dense in $P_{\alpha}\mathcal{O}$ to a single orbit.
Let $v$ be some dense $B$ orbit of $X$, which sits above one of the orbits $\mathcal{O}_{1},...,\mathcal{O}_{k}$. By proposition 6.2 $\pi$ is an isomorphism when restricted to $\Gamma^{v}_{F}(X)$, thus $v$ sits above all the closed orbits in $P_{\alpha}\mathcal{O}$ and it sits above exactly one dense orbit in $P_{\alpha}\mathcal{O}$. Therefore if $\mathcal{O}_{l}$ is a closed orbit in $P_{\alpha}\mathcal{O}$ then $v$ is counted once in $\sum_{\pi(\mathcal{O}_{i})=\pi(\mathcal{O}_{l})}m(\mathcal{O}_{i})=m(\mathcal{O}_{l})$. And if $\mathcal{O}_{l}$ is dense in $P_{\alpha}\mathcal{O}$ $v$ is also counted once in $\sum_{\pi(\mathcal{O}_{i})=\pi(\mathcal{O}_{l})}m(\mathcal{O}_{i})$ as it contains only one of the dense orbits. Thus, in any case $v$ is counted once and therefore the some doesn’t depend on $\mathcal{O}_{l}$. ∎
Corollary 6.6.
The dimension $dim\mathcal{H}(\Gamma_{F}(X))$ is larger or equal to the number of full closed Borel orbits in $\Gamma_{\bar{F}}(X)$.
Proof.
This follows immediately from proposition 5.4 and the previous proposition. ∎
Remark 6.7.
There are case where equality doesn’t hold. For example in the case $SL_{2}/\mathbb{G}_{m}$. In this case $dim\mathcal{H}(\Gamma_{F}(X))=3$ but there are only two full closed orbits in $\Gamma_{\bar{F}}(X)$.
As a result we obtain the following:
Theorem 6.8.
Let $F$ be a finite field, assume that both $G$ and $H$ are split over $F$, then $X$ is qausi split if and only if $St$ is $H$ distinguished.
Proof.
The fact that $X$ being distinguished is a necessary condition can be proven like in Section 4. The other direction follows from the above corollary, proposition 5.2 and proposition 3.1. ∎
7 Harmonic functions in the non archimedean local field case
Let $F$ be a non archimedean local field. Let $\mathcal{O}$ be the integer ring of $F$ and let $k$ be its residue field.
In this Section we will construct a non zero harmonic function on $\Gamma^{0}_{aff,F}(X)$ under the assumptions that both $G$ and $H$ are split, $\sigma$ is defined over $\mathcal{O}$ and $X$ is quasi split, excluding some specific cases.
Our construction will proceed in two independent steps. First, like in proposition 5.5 we construct a function $s$, this time on $H$ orbits of chambers in the affine building $\mathcal{B}$ to $\mathbb{Z}[S]$. Here $S$ is the set of full orbits of $l_{\sigma}$ length zero in $\Gamma^{0}_{aff,F}(X)$. Second, we show that under our assumptions there are full orbits of $l_{\sigma}$ length zero, excluding in some specific cases which we analyze separately.
We begin with the first step.
Definition 7.1.
Recall that the vertices of $\Gamma^{0}_{aff,F}(X)$ can be identified with $H$ orbits of chambers in $\mathcal{B}$, and recall the length function $l_{\sigma}$ on them (see definition 5.1 in [Sht24]). Denote by $S$ the set of chambers of $l_{\sigma}$ length zero which are full (with respect to $\tilde{\Delta}$).
Proposition 7.2.
There exists a function $s:\Gamma^{0}_{aff,F}(X)\rightarrow\mathbb{Z}[S]$ such that:
  1. 1.
    For any $x\in S$ we have $s(x)=x$.
  2. 2.
    For any $E$, a hyper edge of $\Gamma^{0}_{aff,F}(X)$, $\sum_{v\in E}s(v)=0$.
Proof.
We construct $s$ by induction on the $l_{\sigma}$ of a chamber. On full chambers of $l_{\sigma}$ zero it is defined by the first property, on non full chambers of $l_{\sigma}$ zero we define $s$ to be zero.
Now let $\mathcal{C}$ be some chamber in $\mathcal{B}$ and let $\alpha,\beta\in\tilde{\Delta}$ be two simple reflections such that each of the hyper edges labeled $\alpha,\beta$ that contains $\mathcal{C}$, does not contain a chamber whose $l_{\sigma}$ is larger than $l_{\sigma}(\mathcal{C})$. Like in the proof of proposition 5.5 it is enough to verify the result for the maximal sub hyper graph that contains $H\mathcal{C}$ and has hyper edges only with labels $\alpha$ or $\beta$. We denote this hyper graph by $\Gamma_{\alpha,\beta}$.
Each of $\alpha,\beta$ correspond to a codimesnion 1 face of $\mathcal{C}$, denote them by $f_{\alpha}$ and $f_{\beta}$ respectively. Consider $f_{\alpha,\beta}=f_{\alpha}\cap f_{\beta}$. Let $W_{\alpha,\beta}$ be the set of all walls that contain $f_{\alpha,\beta}$.
There a few options.
  1. 1.
    $f_{\alpha,\beta}=\varnothing$.
  2. 2.
    $W_{\alpha,\beta}$ does not contain $\sigma$ stable walls.
  3. 3.
    $W_{\alpha,\beta}$ does contain a $\sigma$ stable wall but $f_{\alpha,\beta}$ does not span a $\sigma$ stable hyper plane.
  4. 4.
    $W_{\alpha,\beta}$ does contain a $\sigma$ stable wall and $f_{\alpha,\beta}$ spans a $\sigma$ stable hyper plane.
We will deal with each case separately. In any case we may choose $\mathcal{A}$, a $\sigma$ stable apartment that contains $\mathcal{C}$.
  1. 1.
    This case may happen only if $f_{\alpha}$ and $f_{\beta}$ are parallel. Let $\mathcal{C}_{\alpha}$ ($\mathcal{C}_{\beta}$) be the chamber in $\mathcal{A}$ that contains $f_{\alpha}$ ($f_{\beta}$) and is different from $\mathcal{C}$. By our assumption $l_{\sigma}(\mathcal{C}_{\alpha})\leq l_{\sigma}(\mathcal{C})$ and also $l_{\sigma}(\mathcal{C}_{\beta})\leq l_{\sigma}(\mathcal{C})$. If we would have a sharp inequity for one of the chambers, it will contradict the other inequity, because if moving in one direction decreases $l_{\sigma}$ then moving in the other direction must increase it. Thus we must have $l_{\sigma}(\mathcal{C})=l_{\sigma}(\mathcal{C}_{\alpha})l_{\sigma}(\mathcal{C}_{\beta})$. This argument works for any choice of $\mathcal{A}$, which means that there are no vertices in $\Gamma_{\alpha,\beta}$ which are connected to $\mathcal{C}$ and have $l_{\sigma}$ less than $l_{\sigma}(\mathcal{C})$. We claim that $\Gamma_{\alpha,\beta}$ must contain only one vertex. Otherwise, by remark 7.9 of [Sht24], $\mathcal{C}$ and $\sigma(\mathcal{C})$ must be on different sides of $f_{\alpha}$ and of $f_{\beta}$ which is clearly impossible.
  2. 2.
    Let $\mathcal{C}^{\prime}\subset\mathcal{A}$ be the reflection of $\mathcal{C}$ across $f_{\alpha,\beta}$. There are no $\sigma$ stable walls that contain $f_{\alpha,\beta}$ so the $\Gamma_{\alpha,\beta}$ is a graph (see Lemma 5.5 of [Sht24]). This graph has a unique vertex with minimal $l_{\sigma}$, this vertex correspondence to $\mathcal{C}^{\prime}$. Thus we can define $s$ on the vertices of this hyper graph by $s(v)=(-1)^{\frac{l_{\sigma}(v)-l_{\sigma}(\mathcal{C}^{\prime})}{2}}\mathcal{C}^{\prime}$. This is well defined again by Lemma 5.5 of [Sht24].
  3. 3.
    Notice, that in this case $W_{\alpha,\beta}$ can contain only a single wall of $\mathcal{A}$ which is $\sigma$ stable. Otherwise, their intersection will be an affine hyper plane spanned by $f_{\alpha,\beta}$.
    First, we claim that one of $f_{\alpha},f_{\beta}$ must lay in the unique $\sigma$ stable wall of $\mathcal{A}$ in $W_{\alpha,\beta}$. Assume the contrary. For any point $x$ in the interior of $\mathcal{C}$, which is close enough to $f_{\alpha}$, the line connecting $x$ and $\sigma(x)$ must intersect $f_{\alpha}$, this is because moving across $f_{\alpha}$ decreases $l_{\sigma}$. The same is true for $f_{\beta}$.
    Thus, by considering points in $\mathcal{C}$ close enough to $f_{\alpha,\beta}$ we can find a codimension one hyper plane that pass through $f_{\alpha,\beta}$ and intersects the interior of $\mathcal{C}$. This hyper plane isn’t a wall of $\mathcal{A}$ and thus is not in $W_{\alpha,\beta}$.
    Intersecting this hyper plane with the $\sigma$ stable wall in $W_{\alpha,\beta}$ shows that $f_{\alpha,\beta}$ spans a $\sigma$ stable hyper plane. This is a contradiction.
    We assume without loss of generality that $f_{\alpha}$ is contained in a $\sigma$ stable wall. Let $E_{\mathcal{C},\alpha}$ be the hyper edge labeled by $\alpha$ which contains $\mathcal{C}$. There are two options, either $E_{\mathcal{C},\alpha}$ contains some chamber whose $l_{\sigma}$ length is smaller than $\mathcal{C}$ or it doesn’t.
    Let $\mathcal{C}^{\prime}\subset\mathcal{A}$ be the reflection of $\mathcal{C}$ across $f_{\alpha,\beta}$. Let $E_{\mathcal{C}^{\prime}}$ be the hyper edge that contains $\mathcal{C}$ and is labeled by the simple reflection that correspondence to the face of $\mathcal{C}^{\prime}$ which is in the hyper plane spanned by $f_{\alpha}$.
    we will prove that $E_{\mathcal{C},\alpha}$ contains a chamber whose $l_{\sigma}$ is less than $l_{\sigma}(\mathcal{C})$ if and only if $E_{\mathcal{C}^{\prime}}$ contains a chamber whose $l_{\sigma}$ is less than $l_{\sigma}(\mathcal{C}^{\prime})$. After proving this, the result will follow as it is to write the options for how $\Gamma_{\alpha,\beta}$ looks.
    1. (a)
      If $E_{\mathcal{C},\alpha}$ does not contain a chamber whose $l_{\sigma}$ is less than $l_{\sigma}(\mathcal{C})$ then $\Gamma_{\alpha,\beta}$ is a graph, which is either a line or a rectangle, depending on the number of different orbits in $E_{\mathcal{C},\alpha}$, which is either 1 or 2. A rectangle graph can be seen in figure 4.
    2. (b)
      If $E_{\mathcal{C},\alpha}$ does contain a chamber whose $l_{\sigma}$ is less than $l_{\sigma}(\mathcal{C})$ then there are four options for how $\Gamma_{\alpha,\beta}$ looks like, according to the elements of $E_{\mathcal{C},\alpha}$. By Theorem 7.12 of [Sht24], the hyper edge $E_{\mathcal{C},\alpha}$ contains one or two elements of minimal $l_{\sigma}$ length and one or two elements of maximal $l_{\sigma}$ length. The hyper graph in the case #$\#E_{\mathcal{C},\alpha}=4$ can be seen in figure 5. If $E_{\mathcal{C},\alpha}$ contains only one element of maximal $l_{\sigma}$ length then the graph will contain only one row of vertices in the top face of the hyper graph. Similarly, if $E_{\mathcal{C},\alpha}$ contains only one element of minimal $l_{\sigma}$ length then the graph will contain only one row of vertices in the bottom face of the hyper graph.
[…]
    Figure 4: A rectangle graph. The red edges are correspond to one simple root and the blue edges to the other.
[…]
    Figure 5: Box hyper graph, the blue edges correspond to one simple root and the red edges to the other. The four vertices connected by red edges on one side of the box are in the same hyper edge. Also the vertices on the other side of the box connected by blue edges are in the same hyper edge. All other hyper edges ar of size 2.
    Now we will prove the assertion above, $E_{\mathcal{C},\alpha}$ contains a chamber whose $l_{\sigma}$ is less than $l_{\sigma}(\mathcal{C})$ if and only if $E_{\mathcal{C}^{\prime}}$ contains a chamber whose $l_{\sigma}$ is less than $l_{\sigma}(\mathcal{C}^{\prime})$.
    The roles of $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are essentially symmetric, so we may assume that $E_{\mathcal{C},\alpha}$ contains a chamber whose $l_{\sigma}$ is less than $l_{\sigma}(\mathcal{C})$ and show the same for $E_{\mathcal{C}^{\prime}}$. Let $\mathcal{C}_{m}\in E_{\mathcal{C},\alpha}$ be such that $l_{\sigma}(\mathcal{C}_{m})<l_{\sigma}(\mathcal{C})$. Let $\mathcal{A}^{\prime}$ be a $\sigma$ stable apartment that contains $\mathcal{C}_{m}$. Let $\mathcal{C}_{m}^{\prime}\subset\mathcal{A}^{\prime}$ be the reflection of $\mathcal{C}_{m}$ across $f_{\alpha,\beta}$. We have $\mathcal{C}^{\prime}_{m}\in E_{\mathcal{C}^{\prime}}$ and $l_{\sigma}(\mathcal{C}^{\prime}_{m})<l_{\sigma}(\mathcal{C}^{\prime})$ because by remark 7.9 of [Sht24] and the fact that $\sigma$ fixes the half spaces of $\mathcal{A}^{\prime}$ separated by the hyper plane spanned by $f_{\alpha,\beta}$, this can be seen again from remark 7.9 of [Sht24] applied to $\mathcal{C}_{m}$ and $\mathcal{C}$.
  4. 4.
    In this case, $\Gamma_{\alpha,\beta}$, is of the form $\Gamma_{k}(X^{\prime})$ for some symmetric space $X^{\prime}$ over a rank 2 group $G^{\prime}$ (see Section 6 of [Sht24]). We know that $\Gamma_{\bar{k}}(X^{\prime})$ is one of the hyper graphs discussed in the proof of proposition 5.5.
    The hyper graph $\Gamma_{k}(X^{\prime})$ can be obtained from on of this graphs by one of the following operations:
    1. (a)
      Delete all the vertices of lower dimension in a hyper edge of type either $N$ or $T$. Also delete all vertices below a deleted vertex, and delete all vertices connected to a deleted vertex by a hyper edge of type $U$ or a type $N$, in the case of type $N$ the condition applies only if the vertex has the same dimension as the deleted vertex connected to it. This correspond to non split tori in proposition 7.5 of [Sht24].
    2. (b)
      Double a vertex of maximal dimension in a hyper edge of type $N$ or $T$. Also double all vertices above a doubled vertex, and double all vertices connected to a doubled vertex by a hyper edge of type $U$. This correspond to split tori in proposition 7.5 of [Sht24].
    It is easy to check that the proposition holds for all the hyper graphs obtained by deleting vertices from the hyper graphs of discussed in the proof of proposition 5.5, any deletion will delete all closed vertices and there is nothing to prove. For the doubling operation, it can also be checked directly. Alternatively, one might use the the argument given in the proof of proposition 6.5.
Remark 7.3.
A simplified version of the arguments given in cases 2 and 3 above, can be used to show that in the proof of proposition 5.5 we may consider only symmetric subgroups of groups of semi simple rank 2.
In order to construct a non zero harmonic function, we only need to show that $S\neq\varnothing$. Before we discuses sufficient conditions for that, we need the following proposition.
Proposition 7.4.
Let $\mathcal{C}$ be a $\sigma$ stable chamber inside a $\sigma$ stable apartment $\mathcal{A}$. Let $T$ be the maximal split torus that corresponds to $\mathcal{A}$, and let $(u_{\alpha})$ be a realization of $G$ defined over $F$ (see Lemma 6.6 of [Hel00]). Let $f$ be a face of $\mathcal{C}$ which lays in a wall defined by a $\sigma$ stable root $\alpha$. The hyper edge of $\Gamma^{0}_{aff,F}(X)$ which consists of $H$ orbits that contain $f$ is of size 1 if and only if we have $\sigma(u_{\alpha}(x))=u_{\alpha}(x)$ for all $x\in F$.
Proof.
For the first direction, assume that $\sigma(u_{\alpha}(x))=u_{\alpha}(x)$ for all $x\in F$, it also implies that $\sigma(u_{-\alpha(x)})=u_{-\alpha(x)}$ for all $x\in F$. Using the group generated by the elements of the form $u_{\pm\alpha(x)}$, the chamber $\mathcal{C}$ can be moved to any other chamber that contains $f$, thus they are all in the same $H$ orbit.
For the second direction, we use the assumption that $\alpha$ is $\sigma$ stable. We have either have, for all $x\in F$ $\sigma(u_{\alpha}(x))=u_{\alpha}(-x)$ or we have, for all $x\in F$, $\sigma(u_{\alpha}(x))=u_{\alpha}(x)$. We want to show that the second case happens.
Assume by contradiction that for all $x\in F$ $\sigma(u_{\alpha}(x))=u_{\alpha}(-x)$. Let $x\in F$ be such that $u_{\alpha}(x)$ fixes $f$, by assumption there is $h\in H$ such that $u_{\alpha}(x)\mathcal{C}=h\mathcal{C}$, thus $u_{\alpha}(-x)\mathcal{C}=\sigma(u_{\alpha}(x)\mathcal{C})=h\mathcal{C}=u_{\alpha}(x)\mathcal{C}$. Thus $u_{\alpha}(2x)$ fixes $\mathcal{C}$, even thou $u_{\alpha}(2x)$ does not fix $\mathcal{C}$. This implies that the valuation of $2x$ is different from the valuation of $x$ which is a contradiction. ∎
Now, we come back and give a sufficient condition for $S\neq\varnothing$.
Proposition 7.5.
Let $G$ be a simple group not of type $A_{2n}$, $\sigma$ an algebraic involution defined over $\mathcal{O}$ and assume that $H=G^{\sigma}$ is split over $\mathcal{O}$. Assume that $X=G/H$ is quasi-split. Then there is a chamber $\mathcal{C}$ in $\mathcal{B}$ with $l_{\sigma}(\mathcal{C})=0$ such that the corresponding vertex in $\Gamma^{0}_{aff,F}(X)$ is full.
Proof.
Consider $o$, the hyperspecial vertex of $\mathcal{B}$ that corresponds to $\mathbf{G}(\mathcal{O})$. The vertex $o$ is $\sigma$ stable because $\sigma$ is defined over $\mathcal{O}$. The $\mathbf{H}(\mathcal{O})$ orbits on the chambers that contain $o$ correspond to $\mathbf{H}(k)$ orbits on the flag variety of $\mathbf{G}(k)$. By proposition 5.2 over $\bar{k}$ there is a closed Borel orbit which is $\Delta$ full. By Theorem 6.1 such an orbit exists also over $k$. This closed Borel orbit corresponds to a chamber $\mathcal{C}\subset\mathcal{B}$ that contains $o$ and is $\sigma$ stable, meaning $l_{\sigma}(\mathcal{C})=0$. This chamber is $\Delta$ full also in $\Gamma^{0}_{aff,F}(X)$, we only need to show that $\mathcal{C}$ is also $\tilde{\Delta}$ full.
Let $T$ be a $\sigma$ stable maximal torus which correspond to a $\sigma$ apartment that contains $\mathcal{C}$. We may choose $\Delta$ to be a set of simple roots which are all positive on $\mathcal{C}$. Let $\gamma$ be the highest root with respect to this choice.
Let $(u_{\alpha})$ be a realization of $G$ defined over $F$ (see Lemma 6.6 of [Hel00]).
The set $\tilde{\Delta}\backslash\Delta$ consists of a single element which corresponds to the unique vertex in the affine Dynkin diagram of $G$, which is not in the usual Dynkin diagram of $G$. The face of $\mathcal{C}$ which correspondence to $\tilde{\Delta}\backslash\Delta$ is parallel to the walls defined by $\gamma$. The root $\gamma$ is $\sigma$ stable so by proposition 7.4 it is enough to show that for all $x\in F$, $\sigma(u_{\gamma}(x))=u_{\gamma}(-x)$.
The chamber $\mathcal{C}$ is $\sigma$ stable, so $\sigma$ induces a map $\sigma:\Delta\rightarrow\Delta$.
There is an involution $\sigma_{0}:G\rightarrow G$ such that $\sigma_{0}(u_{\alpha}(x))=u_{\sigma(\alpha)}(x)$ for all $\alpha\in\Delta$ and $x\in F$. Furthermore, there is some $t\in T$ such that $\sigma=\sigma_{0}\circ Int(t)$, where $Int(t)$ is conjugation by $t$ (see Section 7 of [Spr87]).
Let $\alpha\in\Delta$ be such that $\beta:=\sigma(\alpha)\neq\alpha$. We have, for any $x\in F$, $\sigma_{0}\circ Int(t)\circ\sigma_{0}\circ Int(t)(u_{\alpha}(x))=u_{\alpha}(x)$, therefore $\beta(t)\cdot\alpha(t)=1$.
On the other hand, if $\sigma(\alpha)=\alpha$, we have for any $x\in F$, $\sigma(u_{\alpha}(x))=u_{\alpha}(\alpha(t)x)$ so we must have $\alpha(t)^{2}=1$. We know that $\mathcal{C}$ is $\Delta$ full, therefore by proposition 7.4 we must have $\alpha(t)=-1$.
We can write $\gamma$ as a sum of simple roots $\gamma=\sum_{\alpha\in\Delta}n_{\alpha}\alpha$. Notice that as $\gamma$ is $\sigma$ stable we have $n_{\alpha}=n_{\sigma(\alpha)}$ for any $\alpha\in\Delta$.
We have
$\gamma(t)=\Pi_{\alpha\in\Delta}\alpha(t)^{n_{\alpha}}=\Pi_{\alpha\in\Delta,\sigma(\alpha)=\alpha}\alpha(t)^{n_{\alpha}}=(-1)^{\sum_{\alpha\in\Delta,\sigma(\alpha)=\alpha}n_{\alpha}}=(-1)^{\sum_{\alpha\in\Delta}n_{\alpha}}$
The sum $\sum_{\alpha\in\Delta}n_{\alpha}$ is smaller by one than the Coxeter number of $G$, which is even because $G$ isn’t of type $A_{2n}$, thus $\gamma(t)=-1$.
There are two cases, either $\sigma$ is an inner involution and then $\sigma_{0}$ is trivial or $\sigma$ is an outer involution.
If $\sigma$ is an inner involution, then $\sigma(x_{\gamma}(a))=x_{\gamma}(\gamma(t)a)=x_{\gamma}(-a)$ and we are done.
So we may assume that $\sigma$ is an outer involution. In this case $G$ must be simply laced (see for example [Spr87]). It is enough to show that for any $x\in F$ we have $\sigma_{0}(u_{\gamma}(x))=u_{\gamma}(x)$.
We will prove that for any $\alpha$ a $\sigma$ stable positive root and for all $x\in F$ we have $\sigma_{0}(u_{\gamma}(x))=u_{\gamma}(x)$. We will argue by induction on the weight of the root. We already have this for simple roots by the construction of $\sigma_{0}$. It is enough to show the following:
  1. 1.
    If this property holds for two $\sigma$ stable positive roots $\alpha,\beta$ whose sum is a root, then it also holds for $\alpha+\beta$.
  2. 2.
    If this property holds for for a $\sigma$ stable positive root $\alpha$, then for every simple root $\beta$ such that $\sigma(\beta)\neq\beta$ and such that $\alpha+\beta+\sigma(\beta)$ is a root, the property holds for $\alpha+\beta+\sigma(\beta)$.
By Section 10.2 of [Spr09] there is a bi-linear function $f:X^{*}(T)\times X^{*}(T)\rightarrow\mathbb{Z}$ such that if $\alpha,\beta$ are positive roots whose sum is a root we have for all $x,y\in F$,
$u_{\alpha}(x)u_{\beta}(y)u_{\alpha}(-x)u_{\beta}(-y)=u_{\alpha+\beta}((-1)^{f(\alpha,\beta)}xy)$
Therefore, if $\alpha,\beta$ are positive $\sigma$ stable roots, such thta for all $x\in F$, $\sigma_{0}(u_{\alpha}(x))=u_{\alpha}(x)$ and $\sigma_{0}(u_{\beta}(x))=u_{\beta}(x)$ then the same also holds for $\alpha+\beta$, assuming that $\alpha+\beta$ is a root.
Let $\alpha$ be a positive $\sigma$ stable root such that for all $x\in F$, $\sigma_{0}(u_{\alpha}(x))=u_{\alpha}(x)$. Let $\beta_{1},\beta_{2}$ be two simple roots such that $\sigma(\beta_{1})=\beta_{2}$. Assume that $\alpha+\beta_{1}+\beta_{2}$ is also a root. We will show that for all $x\in F$, $\sigma_{0}(u_{\alpha+\beta_{1}+\beta_{2}}(x))=u_{\alpha+\beta_{1}+\beta_{2}}(x)$ and this will finish the proof.
By the construction of $\sigma_{0}$ we have for all $x\in F$, $\sigma_{0}(u_{\beta_{1}}(x))=u_{\beta_{2}}(x)$.
Let $c\in F^{\times}$ such that for all $x\in F$, $\sigma_{0}(u_{\alpha+\beta_{1}}(x))=u_{\alpha+\beta_{2}}(cx)$. We have for all $x,y\in F$
$\sigma_{0}(u_{\alpha+\beta_{1}}((-1)^{f(\alpha,\beta_{1})}xy))=\sigma_{0}(u_{\alpha}(x)u_{\beta_{1}}(y)u_{\alpha}(-x)u_{\beta_{1}}(-y))=$
$=u_{\alpha}(x)u_{\beta_{2}}(y)u_{\alpha}(-x)u_{\beta_{2}}(-y)=u_{\alpha+\beta_{2}}((-1)^{f(\alpha,\beta_{2})}xy)$
Therefore $c=(-1)^{f(\alpha,\beta_{2})+f(\alpha,\beta_{1})}$. We also have for all $x,y\in F$
$\sigma_{0}(u_{\alpha+\beta_{1}+\beta_{2}}((-1)^{f(\alpha+\beta_{1},\beta_{2})}xy))=\sigma_{0}(u_{\alpha+\beta_{1}}(x)u_{\beta_{2}}(y)u_{\alpha+\beta_{1}}(-x)u_{\beta_{2}}(-y))=$
$=u_{\alpha+\beta_{2}}(cx)u_{\beta_{1}}(u)u_{\alpha+\beta_{2}}(-cx)u_{\beta_{1}}(-y)=u_{\alpha+\beta_{1}+\beta_{2}}((-1)^{f(\alpha+\beta_{2},\beta_{1})}cxy)$
It is enough to show that $(-1)^{f(\alpha+\beta_{1},\beta_{2})}=(-1)^{f(\alpha+\beta_{2},\beta_{1})}c$. This is equivalent to $f(\alpha+\beta_{1},\beta_{2})+f(\alpha+\beta_{2},\beta_{1})+f(\alpha,\beta_{2})+f(\alpha,\beta_{1})$ being even. Using the bi-linearity of $f$, it is enough to show that $f(\beta_{1},\beta_{2})+f(\beta_{2},\beta_{1})$ is even. By 10.2.3 of [Spr09] it is equivalent to $\beta_{1}$ and $\beta_{2}$ being perpendicular. $G$ is not of type $A_{2n}$, so there are no involutions of its Dynkin diagram that send a vertex to a neighboring vertex. Therefore $\beta_{1}$ and $\beta_{2}$ must be perpendicular and we are done. ∎
We are left to deal with the case of $G$ being is of type $A_{2n}$.
Proposition 7.6.
Let $G$ be a simple group split over $F$ of type $A_{2n}$, up to conjugation $G$ has two split symmetric subgroups, $H$ such that $G/H$ is qausi split. Denote by $Z$ the center of $G$, $G/Z\cong PGL_{2n+1}$. The projection of $H$ to $G/Z$ is either conjugate to $P(GL_{n}\times GL_{n+1})$ or to $PO_{2n+1}$. In the first case $S=\varnothing$ and in the second case $S\neq\varnothing$.
Proof.
The classification of quasi split spaces is known (see for example [Spr87]), As we assume that both $G$ and $H$ are split that are no more cases then in the algebraically closed field setting (see [Hel00]).
In the case of $H/Z=P(GL_{n}(F)\times GL_{n+1})(F)$ the involution $\sigma$ is an inner involution, by the same calculation in the proof of the previous proposition we see that any vertex of $\Gamma_{aff,F}(X)$ of $l_{\sigma}$ length zero which $\Delta$ full isn’t $\tilde{\Delta}$ full.
In the case of $H/Z$ being a split form of $PO_{2n+1}(F)$, the involution $\sigma$ is an outer involution. We can identify $X$ with the space of symmetric matrices in $PGL_{2n+1}(F)$. Consider the matrix $g$ which has ones on the anti diagonal and zeros everywhere else. The $I$ orbit of this matrix has $l_{\sigma}$ length zero orbit. It is enough to check that it is full. It is easy to check that for every $s\in\tilde{\Delta}$, $s^{t}gs\neq g$. Thus, the action of $s$ does not fix the orbit and therefore it is full and $S\neq\varnothing$. ∎
We know that the dimension of harmonic functions may be larger than #$\#S$, so a priory $S=\varnothing$ doesn’t promise us that there are no non zero harmonic function on $\Gamma^{0}_{aff,F}(X)$. It is enough to show the the following:
Proposition 7.7.
Let $G=SL_{2n+1}(F)$, $H=GL_{n}(F)\times GL_{n+1}(F)\cap SL_{2n+1}(F)$ and $X=G/H$. The only harmonic function on $\Gamma^{0}_{aff,F}(X)$ is zero.
Proof.
We know that in this case $S=\varnothing$. It is enough to show that in any hyper edge of $\Gamma^{0}_{aff,F}(X)$ there is only one vertex of maximal $l_{\sigma}$ length. After showing this we can prove by induction on $l_{\sigma}$ that any harmonic function must be zero on any chamber. The fact that there are no full vertices with $l_{\sigma}$ length zero is the base of the induction.
Let $\sigma$ be the involution given by conjugation by $-I_{n+1}\times I_{n}\in GL_{n+1}(F)\times GL_{n}(F)$ if $n$ is odd and by $I_{n+1}\times-I_{n}\in GL_{n+1}(F)\times GL_{n}(F)$ if $n$ is even.
By the description of hyper edges given in [Sht24], it is enough to show that for any $\sigma$ stable maximal split torus $T$ of $G$ and for every $\sigma$ stable root $\alpha:T\rightarrow F^{\times}$ of $T$, $\alpha$ restricted to the fixed points of $\sigma$, $T^{\sigma}$, is onto $\alpha:T^{\sigma}\rightarrow F^{\times}$.
All maximal split tori of $G$ are conjugate. We may switch $\sigma$ so some other inner involution and assume that the torus $T$ is the torus of diagonal elements in $G$. Let $w\in N_{G}(T)$ be the element defining the inner involution. We denote the entries of $T$ by $t=(t_{1},...,t_{2n+1})$. Write $w^{-1}tw=(t^{\prime}_{1},...,t^{\prime}_{2n+1})$. Conjugation by $w$ is an involution and therefore for every $11\leq i\leq 2n+$ 1 we have either $t^{\prime}_{i}=t_{i}$ or there is some $j\neq i$ and $c\in F^{\times}$ such that $t^{\prime}_{i}=ct_{j}$ and $t^{\prime}_{j}=c^{-1}t_{i}$.
We have some $\sigma$ stable root $\alpha$, recall all roots are of the form $\alpha_{i,j}(t_{1},...,t_{2n+1})\rightarrow t_{i}/t_{j}$. The root $\alpha$ is $\sigma$ stable so we must have $t^{\prime}_{i}=t_{i}$ and $t^{\prime}_{j}=t_{j}$. There has to be some $k\neq i,j$ such that $t^{\prime}_{k}=t_{k}$ because the number of indices switched by $w$ is even. For any $x\in F^{\times}$ take $t\in T$ with $t_{i}=x,t_{k}=x^{-1}$ and all the rest are 1. The element $t$ is fixed by the involution and $\alpha_{i,j}(t)=x$ so it is onto $F^{\times}$ which is what we wanted to prove. ∎
Now we can deduce the following theorem:
Theorem 7.8.
Let $G$ be a reductive group split over $F$, let $H=G^{\sigma}$, and assume that $H$ is also split over $F$. Let $\mathbf{Z}=Z(\mathbf{G})$ be the center of $\mathbf{G}$. $\mathbf{G}/\mathbf{Z}$ can be written as a product of adjoint simple groups $\mathbf{G}/\mathbf{Z}=G_{1}\times...\times G_{m}$. Then there is a non zero harmonic function on $\Gamma^{0}_{aff,F}(X)$ if and only if the following conditions hold:
  1. 1.
    $X=G/H$ is quasi split.
  2. 2.
    There is no $\sigma$ invariant $G_{i}$, such that $G_{i}\cong PGL_{2n+1}$ and $G^{\sigma}_{i}\cong P(GL_{n}\times GL_{n+1})$.
Proof.
For simple groups we already proved this. Denote $G^{\prime}=G/Z$ and $H^{\prime}=H/Z$. We have, $\Gamma^{0}_{aff,F}(X)=\Gamma^{0}_{aff,F}(G^{\prime}/H^{\prime})$. $\Gamma_{aff,F}(G^{\prime}/H^{\prime})$ can be written as a product of hyper graphs of simple groups or of spaces of the from $K\times K/\Delta K$. Existing of a non zero harmonic function on a product of hyper graphs is equivalent to the existing of a non zero harmonic function on each of them. Thus the results follows from the case of a simple group and the group case.
In the group case it is easy to construct a non zero harmonic function. ∎
Now we deduce the following:
Corollary 7.9.
Let $F$ be a non archimedean local field. Let $G$ and $H$ be split over $F$, let $St_{\chi_{0}}$ be the Steinberg representation of $G$ that corresponds to the character $\chi_{0}$. $St_{\chi_{0}}$ is $H$ distinguished if and only if the conditions of the previous theorem hold.
Proof.
This follows from the previous theorem and proposition 3.2. ∎
8 Steinberg representation of a general unramified character
In the previous section we answered the question when $St_{\chi_{0}}$ is $H$ distinguished. In this section we will answer the question when $St_{\chi}$ is $H$ distinguished for a general unramified character $\chi:G\rightarrow\mathbb{C}^{\times}$.
Recall proposition 3.3. We need a way to check when for a harmonic function $\phi$ on $\Gamma^{1}_{aff,F}(X)$ the condition: for any chamber $\mathcal{C}$ and $h\in H$, $\phi(H^{1}h\mathcal{C})=\phi(H^{1}\mathcal{C})\chi^{-1}\chi_{0}(\omega(h^{-1}))$. Here $H^{1}=H\cap Ker(\omega)$ is a normal subgroup of $H$.
We will construct a subgroup of $\Omega$ and show that to check this condition is the same as checking that $\chi^{-1}\chi_{0}$ is trivial on this subgroup.
We will need the following results.
Proposition 8.1.
  1. 1.
    Every $\sigma$ stable apartment that contains a chamber of $l_{\sigma}$ length 0 corresponds to a maximal split tours $T$ such that $T\cap H$ is a maximal torus of $H$.
  2. 2.
    Every two maximal split tori $T_{1},T_{2}$ such that $T_{1}\cap H$ and $T_{2}\cap H$ are maximal tori of $H$ are $H$ conjugate.
Proof.
  1. 1.
    Let $\mathcal{A}$ be a $\sigma$ stable apartment and let $T$ be the maximal split torus that corresponds to $\mathcal{A}$. Assume that $\mathcal{C}\subset\mathcal{A}$ is a chamber such that $l_{\sigma}(\mathcal{C})=0$. Let $T^{+}=T\cap H$, assume by contradiction that $T^{+}$ is not a maximal torus of $H$, let $M=C_{G}(T^{+})$ be the centralizer of $T^{+}$. The group $M$ is a Levi of $G$ on which $\sigma$ acts. Either $M$ contains a $\sigma$ invariant root subgroup or all $\sigma$ stable tori of $M$ are $H$ conjugate. We know that the second option doesn’t happen because $T^{+}$ isn’t a maximal torus of $G$. Therefore there is a root $\alpha$ such that its root subgroup $U_{\alpha}$ is contained in $M$ and is $\sigma$ invariant.
    Let $W\subset\mathcal{A}$ be a wall on which $\alpha$ is constant. The wall $W$ is $\sigma$ stable, $\sigma(\mathcal{C})=\mathcal{C}$ so $\sigma$ must preserve the half planes defined by $W$. The translation defined by $\alpha^{\vee}$, which is the translation perpendicular to $W$, commutes with $\sigma$. By proposition 5.7 of [Sht24], we may assume that this translation is in $H\cap T=T^{+}$, which contradicts $U_{\alpha}\subset C_{G}(T^{+})$.
  2. 2.
    Let $T_{1},T_{2}$ be two maximal split tori of $G$, such that $T_{1}\cap H$, $T_{2}\cap H$ are maximal tori of $H$. The tori $T_{1}\cap H$ and $T_{2}\cap H$ are $H$ conjugate, we may assume that $T_{1}\cap H=T_{2}\cap H$. Denote $T^{+}=T_{1}\cap H=T_{2}\cap H$. Consider $M=C_{G}(T^{+})$, $\sigma$ acts on $M$ and every fixed point of $\sigma$ on $M$ is an element of $C_{H}(T^{+})=T^{+}$. Clearly $T_{1},T_{2}\subset M$. Any maximal split torus $T\subset M$ contains $T^{+}$ and thus its $\sigma$ split part $\{t\in T|\sigma(t)=t^{-1}\}$ is a maximal $\sigma$ split torus of $M$. By 10.3 of [HW93] and the fact that $M$ is split we get that the $\sigma$ split parts of $T_{1}$ and $T_{2}$ are $T^{+}$ conjugate. This can only happen if $T_{1}=T_{2}$.
In the rest of the section we assume that both $G$ and $H$ are split over $\mathcal{O}$. Let 00 be the hyper special vertex of $\mathcal{B}$ that correspond to $G(\mathcal{O})$. Let $\mathcal{A}$ be the $\sigma$ stable apartment of minimal rank that contains 00.
Proposition 8.2.
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be two $\sigma$ stable chambers in $\mathcal{A}$, let $h\in H$ such that $h\mathcal{C}_{1}=\mathcal{C}_{1}$ then there is $h^{\prime}\in H$ such that $h^{\prime}\mathcal{C}_{2}=\mathcal{C}_{2}$ and $\omega(h^{\prime})=\omega(h)$.
Proof.
Let $T$ be the $\sigma$ stable maximal split torus that correspondence to $\mathcal{A}$. We can lift the Weyl group $N_{G}(T)/T$ to a subgroup of $N_{G}(T)$ of element that fix $0\in\mathcal{A}$, denote this subgroup by $W$.
Notice that any $\sigma$ stable chamber in $\mathcal{A}$ is in the $H$ orbit of a chamber that contains 00. Thus, we may find $w_{2}\in W$ such that $w_{2}\mathcal{C}_{1}=\mathcal{C}_{2}$. Applying $\sigma$ we get that also $\sigma(w_{2})\mathcal{C}_{1}=\mathcal{C}_{2}$, therefore we can find $t_{2}\in T^{0}$ such that $\sigma(w_{2})=t_{2}w_{2}$. It implies $\sigma(t_{2})=t^{-1}_{2}$, we can further assume that $t_{2}$ is $1$ modulo the uniformaizer and by Hensel’s lemma it has a square root in $T^{-}$, denote it by $t^{\frac{1}{2}}_{2}$. Consider $h^{\prime}=w^{-1}_{2}t^{-\frac{1}{2}}_{2}ht^{\frac{1}{2}}_{2}w_{2}$, we claim that $h^{\prime}\in H$. Computing we get that $\sigma(h^{\prime})=w^{-1}_{2}t^{-1}_{2}t^{\frac{1}{2}}_{2}ht^{-\frac{1}{2}}_{2}t_{2}w_{2}=h^{\prime}$. Clearly $\omega(h^{\prime})=\omega(h)$ and $h^{\prime}\mathcal{C}_{2}=\mathcal{C}_{2}$ so we are done. ∎
Now we can prove:
Proposition 8.3.
Let $\mathcal{C}$ be a chamber such that $l_{\sigma}(\mathcal{C})=0$, then $H_{\mathcal{C}}=\{\omega(h)|h\in H,h\mathcal{C}=\mathcal{C}\}\subset\Omega$ doesn’t depend on $\mathcal{C}$.
Proof.
By proposition 8.1 we may assume that $\mathcal{C}\subset\mathcal{A}$. Then, by proposition 8.2 we are done. ∎
Recall we care about harmonic functions on $\Gamma_{aff,F}^{1}(X)$. By the same argument as in the proof of proposition 7.2, harmonic functions on $\Gamma_{aff,F}^{1}(X)$ can be constructed by specifying their values on the $H^{1}$ orbits on chambers of $l_{\sigma}$ length 00 and extending using a function $s$ similar to the one constructed in proposition 7.2. There are finitely many $H^{1}$ orbits of chambers of $l_{\sigma}$ length 00. Denote the set of these orbits by $S$. Clearly $H$ acts on $S$.
Proposition 8.4.
Assume that $S\neq\varnothing$ and that there is no chamber $\mathcal{C}$ with $l_{\sigma}(\mathcal{C})=0$ and $h\in H$ with $h\mathcal{C}=\mathcal{C}$ and $\chi^{-1}\chi_{0}(\omega(h))\neq 1$ then there is a non zero function $\phi\in\mathcal{H}(\Gamma^{1}_{aff,F}(X))$ such that for any chamber $\mathcal{C}$ and $h\in H$, $\phi(H^{1}h\mathcal{C})=\phi(H^{1}\mathcal{C})\chi^{-1}\chi_{0}(\omega(h^{-1}))$.
Proof.
We have a function $s:\Gamma^{1}_{aff,F}(X)\rightarrow\mathbb{Z}[S]$ like the one constructed in the proof of proposition 7.2. We define $\phi$ on $S$ and then extend it to $\Gamma^{1}_{aff,F}(X)$. Choose some $x\in S$, we define $\phi(hx)=\chi^{-1}\chi_{0}(\omega(h^{-1}))$ for elements in the $H$ orbit of $x$. We define $\phi$ to be zero on elements of $S$ not in the $H$ orbit of $x$. $\phi$ is well defined on $S$ by our assumptions.
Now extend $\phi$ to $\mathbb{Z}[S]$ and to $\Gamma^{1}_{aff,F}(X)$ by the function $s$.
Clearly $\phi$ is harmonic and non zero. We just have to check that for every $h\in H$ and every chamber $\mathcal{C}$, $\phi(H^{1}h\mathcal{C})=\phi(H^{1}\mathcal{C})\chi^{-1}\chi_{0}(\omega(h^{-1}))$.
Notice that the action of $H$ on the vertices of $\Gamma^{1}_{aff,F}(X)$ commutes with $s$ and the action of $H$ on $\mathbb{Z}[S]$.
We have $s(H^{1}h\mathcal{C})=s(H^{1}\mathcal{C})$. By the construction of $\phi$ and its values on $S$ we have $\phi(H^{1}h\mathcal{C})=\phi(hs(H^{1}\mathcal{C}))=\chi^{-1}\chi_{0}(\omega(h^{-1}))\phi(s(H^{1}\mathcal{C}))$. ∎
Now we deduce:
Theorem 8.5.
Let $G$ be a reductive group split over $\mathcal{O}$, let $H=G^{\sigma}$, and assume that $H$ is split over $F$. Let $\mathbf{Z}=Z(\mathbf{G})$ be the center of $\mathbf{G}$. $\mathbf{G}/\mathbf{Z}$ can be written as a product of adjoint simple groups $\mathbf{G}/\mathbf{Z}=G_{1}\times...\times G_{m}$.
Let $\chi:G\rightarrow\mathbb{C}^{\times}$ be an unramified character, and let $St_{\chi}$ be the corresponding Steinberg representation. $St_{\chi}$ is $H$-distinguished if the following conditions hold:
  1. 1.
    $X$ is quasi-split.
  2. 2.
    There is no $\sigma$ invariant $G_{i}$, such that $G_{i}\cong PGL_{2n+1}$ and $G^{\sigma}_{i}\cong P(GL_{n}\times GL_{n+1})$.
  3. 3.
    Let $\mathcal{C}$ be a chamber such that $l_{\sigma}(\mathcal{C})=0$, $\chi^{-1}\chi_{0}$ is trivial on $\Omega_{H}=\{h\in H|h\mathcal{C}=\mathcal{C}\}$.
9 Interpretation of the main results in terms of the dual group
Let $G^{\vee}$ be the Langlands dual of $G$, there are a few different definitions for $G_{X}^{\vee}$, the dual group of $X$. We follow the one given by Takeda for symmetric spaces in [Tak23]. In [Tak23] it is stated without a proof that their definition for $G_{X}^{\vee}$ is equivalent to the definition given in [GN10]. Notably, this definition is different from the one given in [KS17]. In [Tak23] a map $\iota:SL_{2}\times G_{X}^{\vee}\rightarrow G^{\vee}$ is defined, it is an embedding of $G^{\vee}_{X}$ into $G^{\vee}$. In this section we will study when the Langlands parameter of the Steinberg representation factors through $\iota$.
Let $\mathcal{U}$ be the variety of unipotent elements of $G^{\vee}$ and let $\mathcal{U}_{X}$ be the variety of unipotent elements of $G^{\vee}_{X}$.
Let $W_{F}$ be the Weil-Deligne group of $F$. Let $\phi_{St}:W_{F}\times SL_{2}(\mathbb{C})\rightarrow G^{\vee}$ be the expected parameter of the Steinberg representation under the expected local Langlands correspondence (see for example section 1.4 of [CS24]).
The conjectures of [SV17] (see also conjecture 1.1 of [Tak23]) imply that the distinction of the Steinberg representation should be detected by the factoring of $\phi_{St}$ through $\iota$. In this section we prove the following:
Theorem 9.1.
Assume that $H$ is split over $F$. $St_{\chi_{0}}$ is $H$ distinguished if and only if $\phi_{St}$ factors through $\iota$.
Remark 9.2.
Notice that in Theorem 9.1, the parameter is the parameter of $St=St_{1}$ and not of $St_{\chi_{0}}$
Before we begin the proof we will recall some properties of $G^{\vee}_{X}$ and $\iota$ from [Tak23].
9.1 Properties of $G^{\vee}_{X}$
Let $T$ be a maximal $\sigma$ stable split torus of $G$ such that $T^{-}=\{t\in T|\sigma(t)=t^{-1}\}^{0}$ is a maximal split torus on which $\sigma$ acts as inverse.
Let $R$ be the set of roots of $G$ with respect to $T$ and let $\Delta$ be a set of simple roots of $R$, chosen such that if $\alpha$ is a positive root such that $\sigma(\alpha)\neq\alpha$ then $\sigma(\alpha)$ is negative (see 3.4 of [Tak23]).
Proposition 9.3 (See 7.1 of [Tak23]).
Let $2\rho_{0}=\sum_{\alpha\in\Delta,\sigma(\alpha)=\alpha}\alpha$.
The image of $\iota$ on the torus of $SL_{2}(\mathbb{C})$ is the image of the $G^{\vee}$ coroot $2\rho_{0}$.
Proposition 9.4 (See Lemma 3.2 of [Tak23]).
Assume that the root system of $G$ is irreducible. Let $w_{0}$ be the longest element of the Weyl group. Consider $\sigma^{*}:R\rightarrow R$ defiend by $\sigma^{*}(\alpha)=-\sigma(w_{0}(\alpha))$. Then $\sigma^{*}$ is an automorphism of the Dynkin diagram of $G$ such that for every $\alpha\in\Delta\backslash\Delta_{0}$ we have $\sigma(\alpha)+\sigma^{*}(\alpha)\in\mathbb{Z}\Delta_{0}$
Definition 9.5 (See 4.1 of [Tak23]).
Let $\alpha\in\Delta\backslash\Delta_{0}$, define $\Phi_{\alpha}=\mathbb{Z}(\Delta_{0}\cap\{\alpha,\sigma(\alpha)\})\cap R$.
Define the Dynkin diagram of $\alpha$ to be the Dynkin diagram of the irreducible component of $\Phi_{\alpha}$ that contains $\alpha$. We consider the Dynkin diagram with roots in $\Delta_{0}$ colored black and the action of $\sigma^{*}$ drawn on the diagram.
These diagrams were classified in [Tak23] and appear in table 1 of [Tak23].
We will interested in diagrams with no black vertices. Such diagrams from table 1 of [Tak23] appear in Table 1, the numbering is taken from [Tak23].
Table 1: Rank one involutions with no black vertices
No. $\Phi$ Diagram Type
1 $A_{1}\times A_{1}$dynkin[labels=,, involutions=1¡above¿[σ^*]2, edge/.style=white]Aoo 3
2 $A_{1}$dynkinAo 2
5 $A_{2}$dynkin[involutions=1¡above¿[σ^*]2]Aoo 1
There are four options for each simple root $\alpha\in\Delta$ (see 4.3 of [Tak23]):
  • $\sigma(\alpha)=\alpha$. Denote the set of this simple roots by $\Delta_{0}$.
  • $\alpha-\sigma(\alpha)$ is a root. Those are called roots of type 1.
  • $\frac{1}{2}(\alpha-\sigma(\alpha))$ is a root. Those are called roots of type 2.
  • None of the above holds. Those are called roots of type 3.
Somethings can be said for roots of type 3. For a root $\alpha$ of type 3 there are exist a unique pair of roots $\alpha_{1},\alpha_{2}$ characterized by proposition 5.2 of [Tak23], one of the conditions is $\alpha-\sigma(\alpha)=\alpha_{1}+\alpha_{2}$.
Proposition 9.6 (See subsection 5.3 of [Tak23]).
Let $\alpha\in\Delta$ such that $\sigma(\alpha)\neq\alpha$. The group $G^{\vee}_{X}$ contains the following root subgroups:
  1. 1.
    $\{u_{\alpha-\sigma(\alpha)}(x)|x\in\mathbb{C}\}$ if $\alpha$ is of type 1.
  2. 2.
    $\{u_{\frac{1}{2}(\alpha-\sigma(\alpha))}(x)|x\in\mathbb{C}\}$ if $\alpha$ is of type 2.
  3. 3.
    $\{u_{\alpha_{1}}(x)u_{\alpha_{2}}(x)|x\in\mathbb{C}\}$ if $\alpha$ is of type 3.
9.2 Proof of Theorem 9.1
We begin with recalling how $\phi_{St}$ is expected to look, see for example Lemma 1.8 in [CS24].
Denote by $\lambda_{St}:W_{F}\rightarrow G^{\vee}$ the infinitesimal character of $St$, $\lambda_{St}(w)=\phi_{St}(w,\begin{pmatrix}|w|^{\frac{1}{2}}&0\\ 0&|w|^{-\frac{1}{2}}\end{pmatrix})$.
Proposition 9.7 (See Lemma 1.8 of [CS24]).
The parameter $\phi_{St}:W_{F}\times SL_{2}(\mathbb{C})\rightarrow G^{\vee}$ satisfies the following:
  1. 1.
    $\phi_{St}$ is trivial on $W_{F}$.
  2. 2.
    $\phi_{St}$ maps $\begin{pmatrix}1&x\\ 0&1\end{pmatrix}$ to the open $G^{\vee}$ orbit of $\mathcal{U}$.
  3. 3.
    The infinitesimal character $\lambda_{St}$ is the same as the infinitesimal character of the trivial representation.
We have a characterization of when $St_{\chi_{0}}$ is distinguished given by Corollary 7.9. We need to check that these conditions are equivalent to $\phi_{St}$ factoring through $\iota$.
Proposition 9.8.
Assume that the parameter $\phi_{St}$ factors through $\iota$, then $X$ is quasi split.
Proof.
let $t\in SL_{2}(\mathbb{C})$ be semisimple, $\iota((t,1))$ is semisimple and commutes with $\phi_{St}(\begin{pmatrix}1&x\\ 0&1\end{pmatrix})$, an element in the open orbit of $\mathcal{U}$. Thus $\iota((t,1))$ must be in the center of $G^{\vee}$. By proposition 9.3 it means that there are not roots on fixed by $\sigma$. This is equivalent to being quasi split as it implies that for the Borel subgroup which corresponds to the choice of simple roots $\Delta$, $B\cap\sigma(B)$ is a torus. ∎
Let $\mathbf{Z}=Z(\mathbf{G})$ be the center of $\mathbf{G}$. $\mathbf{G}/\mathbf{Z}$ can be written as a product of adjoint simple groups $\mathbf{G}/\mathbf{Z}=G_{1}\times...\times G_{m}$.
We need to show that the opposite direction of proposition 9.8 also holds, except in the case, where there is a $\sigma$ invariant $G_{i}$ such that $G_{i}\cong PGL_{2n+1}$ and $G^{\sigma}_{i}\cong P(GL_{n}\times GL_{n+1})$.
Proposition 9.9.
$\iota(\mathcal{U}_{X})$ intersects the open orbit of $\mathcal{U}$ if and only if the following conditions hold:
  1. 1.
    $X=G/H$ is quasi split.
  2. 2.
    There is no $\sigma$ invariant $G_{i}$, such that $G_{i}\cong PGL_{2n+1}$ and $G^{\sigma}_{i}\cong P(GL_{n}\times GL_{n+1})$.
Proof.
We already showed that $X$ being quasi split is a necessary condition. In particular $\Delta_{0}=\varnothing$.
Now we assume the second condition and will show that $\iota(\mathcal{U}_{X})$ intersects the open orbit of $\mathcal{U}$.
Let $\alpha\in\Delta$ be some simple root, its diagram (see definition 9.5) can’t contain black vertices (as $\Delta_{0}=\varnothing$), so it must be one of the diagrams of table 1.
$\iota(\mathcal{U}_{X})$ intersects the open orbit of $\mathcal{U}$ if and only if on the level of Lie algebras $\mathcal{N}_{X}$ contains an element of the form $\sum_{\alpha\in\Delta}n_{\alpha}(x_{\alpha})$. Here, $\mathcal{N}_{X}$ is the nilpotent cone of $G^{\vee}_{X}$ and for every $\alpha\in\Delta$ $n_{\alpha}(x_{\alpha})$ is a non zero element in the root space of $G^{\vee}$ which correspondence to the root $\alpha$.
Now we will go over the three cases of table 1.
  1. 1.
    In the first case $\alpha$ is of type 3 and $\alpha_{1}=\alpha$, $\alpha_{2}=-\sigma(\alpha)$. By proposition 9.6 we have $\{n_{\alpha}(x)+n_{\sigma(\alpha)}(x)|x\in\mathbb{C}\}\subset\mathcal{N}_{X}$.
  2. 2.
    In the second case $\alpha$ is of type 2 and $\sigma(\alpha)=-\alpha$. By proposition 9.6 we have $\{n_{\alpha}(x)|x\in\mathbb{C}\}\subset\mathcal{N}_{X}$.
  3. 3.
    In the third and last case, $\alpha$ is of type 2, $\sigma(\alpha)\neq-\alpha$ and $\alpha,\sigma(\alpha)$ aren’t perpendicular. From the classification of involutions (see for example [Spr87]), this contradicts the second assumption of our propositioning. Therefore this case can’t happen.
Now it is easy to see that $\iota(\mathcal{U}_{X})$ must intersect the open orbit of $\mathcal{U}$.
For the opposite condition, it is enough to check that in the case of $G=PGL_{2n+1}$ and $H=P(GL_{n}\times GL_{n+1})$, $\iota(\mathcal{U}_{X})$ doesn’t intersect the open orbit of $\mathcal{U}$. In this case $G^{\vee}=SL_{2n+1}$ and $G^{\vee}_{X}=Sp_{2n}$. The result is evident. ∎
Now we can deduce Theorem 9.1.
Proof of Theorem 9.1.
For the first direction, assume that $\phi_{St}$ factors through $\iota$, then by proposition 9.7 $\iota(\mathcal{U}_{X})$ intersects the open $G^{\vee}$ orbit of $\mathcal{U}$. By proposition 9.9 and Theorem 8.5, $St_{\chi_{0}}$ is $H$ distinguished.
For the second direction, we assume that $St_{\chi_{0}}$ is $H$ distinguished and we need to show that $\phi_{St}$ factors through $\iota$. By proposition 9.7 it is enough to show that $\phi_{St}:SL_{2}(\mathbb{C})\rightarrow G^{\vee}$ factors through $\iota$.
Recall that we fixed $T$ be a maximal $\sigma$ stable split torus of $G$ such that $T^{-}=\{t\in T|\sigma(t)=t^{-1}\}^{0}$ is a maximal split torus on which $\sigma$ acts as inverse. Let $B$ be a a Borel containing $T$ which correspondence to the positive roots $R^{+}$ we fixed in subsection 9.1. Let $U$ be the unipotent radical of $B$.
Fix $x\in\mathbb{C}$. By our proof proposition 9.9 we can conjugate $\phi_{St}$ and find $u\in G^{\vee}_{X}$ such that $\phi_{St}(\begin{pmatrix}1&x\\ 0&1\end{pmatrix})=\iota(u)\in U$.
The infinitesimal character of $St$ is the same as of the trivial representation. By Theorem 8.2 of [Tak23] the Langlands parameter of the trivial representation factors through $\iota$. Furthermore we can find $t\in G^{\vee}_{X}$ such that $\phi_{St}(\begin{pmatrix}x^{\frac{1}{2}}&0\\ 0&x^{-\frac{1}{2}}\end{pmatrix})=\iota(t)\in T$.
Using Jacobson-Morozov (see for example Section 3 of [BMIY24]) we can extend the map sending $\begin{pmatrix}1&x\\ 0&1\end{pmatrix}$ to $u$ and $\begin{pmatrix}x^{\frac{1}{2}}&0\\ 0&x^{-\frac{1}{2}}\end{pmatrix}$ to $t$, to the entire group $SL_{2}(\mathbb{C})$